### v1

# Dynamical phase transitions in the non-reciprocal Ising model

Abstract
Non-reciprocal interactions in many-body systems lead to time-dependent states, commonly observed in biological, chemical, and ecological systems. The stability of these states in the thermodynamic limit, as well as the criticality of the phase transition from static to time-dependent states remains an open question. To tackle these questions, we study a minimalistic system endowed with non-reciprocal interactions: an Ising model with two spin species having opposing goals. The mean-field equation predicts three stable phases: disordered, ordered, and a time-dependent swap phase. Large scale numerical simulations support the following: (i) in 2D, the swap phase is destabilized by defects; (ii) in 3D, the swap phase is stable, and has the properties of a time-crystal; (iii) the transition from disorder to swap in 3D is characterized by the critical exponents of the 3D XY model, in agreement with the emerging continuous symmetry of time translation invariance; (iv) when the two species have fully anti-symmetric couplings, the static-order phase is unstable in any dimension due to droplet growth; (v) in the general case of asymmetric couplings, static order can be restored by a droplet-capture mechanism preventing the droplets from growing indefinitely. We provide details on the full phase diagram which includes first- and second-order-like phase transitions and study the coarsening dynamics of the swap as well as the static-order phases.
I Introduction
Many-body systems with non-reciprocal interactions emerge in contexts ranging from neuroscience [1, 2, 3] and social networks [4, 5, 6] to ecology [7, 8, 9, 10] and open quantum systems [11, 12, 13]. These systems exhibit rich collective behavior including time-dependent states that can arise because the dynamics does not follow from the minimisation of a single potential function [14, 15, 16, 17, 18, 19, 20, 21]. However, it remains unknown whether these time-dependent states can be thought of as proper non-equilibrium phases of matter, in the sense of the term used in statistical physics; and if so, whether they exhibit well-defined phase transitions and critical behavior. In particular, we are interested in the fate of these time-dependent states in systems made of locally coupled units subject to thermal noise in finite spatial dimensions when they are taken to the thermodynamic limit.
In order to address this question, we consider a non-reciprocal version of the Ising model, in which two species of spins have incompatible goals: spins of species A tend to align with spins of species B, while spins B tend to anti-align with spins A. This leads to a cyclic dynamics where spins of each species tend to flip in turn. We study this system through analytical arguments and extensive Monte-Carlo simulations. A short account of our results is presented in Letter form in Ref. [22]; we summarize here the most salient points. First, numerical simulations indicate that in our model, a time-dependent phase with long-range spatial and temporal order indeed exists in three dimensions. In this phase, the magnetization oscillates for arbitrarily long times. The system acts as a perfect clock in the thermodynamic limit, in the sense that the coherence time of the oscillations diverges with the size of the system. Second, a continuous phase transition between this oscillatory phase and a disordered (i.e. paramagnetic) phase, corresponding to a Hopf bifurcation in mean-field, is observed. Numerical evidence shows that it is associated with scale invariance, and the corresponding static critical exponents are compatible with the 3D XY universality class. Third, we found that a static ordered phase (similar to a ferromagnetic phase) can exist when non-reciprocal interactions are present, but only if there is some symmetric coupling between the species. However, we found no evidence of a direct transition between the oscillatory phase and the static-order phase, in contrast with mean-field predictions.
II Background
II.1 Non-reciprocal interactions between groups
There is a rich literature on nonequilibrium lattice models, including variants of the Ising model with asymmetric couplings exhibiting a variety of complex phenomena [23, 24, 25, 26, 27]. A distinguishing feature of the non-reciprocal Ising model introduced in this paper is the presence of non-reciprocal interactions between two groups of spins of comparable sizes. This is a rather generic feature that arises in a variety of physical contexts, some of which can be captured by Ising or Ising-like minimal models. Examples include:
  • Nonlinear parametric resonators, that can be mapped to Ising spins. Networks of coupled parametric oscillators have been used as platforms to realize (reciprocal) “Ising machines” that solve combinatorial optimization problems [28, 29, 30, 31, 32, 29, 33, 34]. This is essentially because a single of these resonators implements a double-well potential, and a linear coupling effectively produces a (reciprocal) Ising-like coupling. Reference [35] shows experimentally and analytically that driven non-identical parametric resonators can be mapped into non-reciprocally coupled Ising spins. A network of such resonators could directly realize the NR Ising model discussed in the present work, as well as implement models of learning with non-reciprocal couplings between the artifical neurons [36, 3, 2, 37, 38, 39, 40].
  • Models of collective opinion dynamics, such as in the voter model [41, 23], where the values of the spins represent different opinions. Different communities may have asymmetric interactions: for instance, this is the case of conformists and contrarians (or alternatively trend-setters) [42, 43, 41, 44, 45, 46, 47, 48, 49]. Similarly, spin models are used to model game theory and decision making of economic agents [6].
  • Models of the collective response of transmembrane receptors, that have been put forward for the high sensitivity of cells to external chemicals in chemotaxis [50, 51, 52, 53, 54]. It has been argued that non-reciprocal couplings between the receptors and the kinases that serve as an output signal is required to account for experimental observations [55, 56, 57].
  • Models of neurons based on networks of spins, such as the Hopfield model [58, 59], can contain asymmetric interactions [1, 60, 61, 62, 63, 36, 3, 2], that account not only for existence of asymmetric neural connections but also for the presence of different groups of neurons called excitatory and inhibitory, which correspond to different signs of the couplings [64, 36]. This approach has also been extended to larger brain structures like different brain regions: active/non-active regions are represented by spin variables while their interaction is mimicked by the spin alignment interaction [65].
  • Models of spinor Bose-Einstein condensates based on two populations of Ising spins in a cavity (a Dicke model) that effectively interact non-reciprocally have been considered [11, 66, 67] to account for non-stationary dynamics observed in experiments [68].
  • Model inference in which symmetric and asymmetric Ising models have also been used as a variational ansatz [69, 70, 71, 72, 37, 73].
In several of these classes of systems, it has been conjectured that criticality may play a nontrivial role [74, 75, 76, 77, 78, 79, 80]. Our non-reciprocal Ising model can be seen as a stripped down version of these situations. In realistic cases, (i) there may be more than two populations (ii) the spins may take more than two possible states and (iii) the network topology may be more complex than $\mathbb{Z}^{d}$. While these differences are likely to induce additional complexities and phenomena, we expect that the basic features of the non-reciprocal Ising model will persist for systems in the same universality class [23, 75]. In particular, we expect that locally-coupled many-body systems undergoing a Hopf bifurcation in mean-field will enter this universality class even if they are not originally composed of coupled spin-like degrees of freedom.
II.2 Phase transitions to time-dependent states
An underlying motivation for this work is to construct a bridge between bifurcations in non-variational systems and phase transitions out of equilibrium. In the equilibrium Ising model, for instance, the ferromagnet/paramagnet transition corresponds to a pitchfork bifurcation in mean-field. This means that the scalar order parameter $\phi$ (e.g. the magnetization) follows the dynamics $(d/dt)\phi=\alpha\phi-\phi^{3}$ in which $\alpha\propto T-T_{\text{c}}$ is proportional to the distance to the critical temperature. Conversely, the thermal field theory corresponding to locally-coupled pitchfork bifurcations where
$\partial_{t}\phi=\alpha\phi-\phi^{4}+D\Delta\phi+\eta(t,r),$ (1)
in which $\eta$ is a thermal noise, leads upon renormalization to the Ising critical point [27]. This dynamics can be seen as the model A dynamics (in the formalism of Hohenberg and Halperin [81]) associated to a $\phi^{4}$ free energy $\mathcal{F}$, namely
$\Gamma\partial_{t}\phi=-\frac{\delta\mathcal{F}}{\delta\phi}$ (2)
where $\Gamma$ is a kinetic coefficient.
In non-reciprocal and other non-equilibrium systems, the dynamics is typically non-variational,
$\Gamma\partial_{t}\phi=f[\phi]\neq-\frac{\delta\mathcal{F}}{\delta\phi}$ (3)
so bifurcations with no analogue at equilibrium can occur, in particular those leading to limit cycles. The simplest instance is the Hopf bifurcation, which describes the appearance of a limit cycle from a fixed point (unless otherwise specified, “fixed points” in this paper refer to fixed points of the physical dynamics, not fixed points of a renormalization group (RG) flow), see Refs. [82, 83]. This bifurcation is described by the equation
$\dot{z}=(\alpha+i\omega)z-|z|^{2}z$ (4)
for the complex number $z$ and takes place at $\alpha=0$. When $z$ is replaced by a complex field $\psi$ with appropriate diffusive couplings, the corresponding equation is known as the complex Ginzburg-Landau equation [84]. The fate of this bifurcation has been analyzed using perturbative RG in Refs. [85, 86], and later reproduced and generalized in Refs. [87, 88]. In these works, a phase transition generalizing the Hopf bifurcation is predicted based on a $d=4-\epsilon$ expansion and formally related to the RG fixed point of model A dynamics with $O(2)$ symmetry (i.e., to the XY universality class). Using this analogy Refs. [85, 86] predict that this phase transition, which leads to an oscillating phase with long-range order, should exist for space dimension $d>2$, while no coherent oscillations are expected in $d<2$, and quasi-long-range order in conjectured in $d=2$ 111 References [85, 86] conjecture that quasi-long-range may be present at $d=2$, but more recent studies based on a mapping to a version of the Kardar-Parisi-Zhang (KPZ) equation with compact variables suggest that the phase with algebraic order does not persist in the limit of large systems (nor an equivalent of the Berezinskii-Kosterlitz-Thouless (BKT) transition) but are replaced with a stretched exponential decay of correlations corresponding to the rough phase of the KPZ equation [101, 102]. In addition, Refs. [156] argue that the modified effective interaction between spiral defects (compared to that between XY vortices) allows for a finite density of defects at any temperature, potentially suppressing the algebraic order. . To the best of our knowledge, these predictions have only been tested qualitatively, in Refs. [90, 91], where numerical simulations of coupled three-state oscillators were performed for lattices of size $L^{d}$ going up to $L=80$ in $d=3$. A comparison of the rescaled finite-size scaling data of Refs. [90, 91] with Ising and XY exponents suggested a better collapse with XY exponents. The results of our large-scale simulations with lattices of sizes up to $L=320$ in $d=3$ are in agreement with the predictions of Refs. [85, 86] and with the qualitative observations of Refs. [90, 91]. In particular, we obtain values of the critical exponents that are compatible with XY exponents but not with Ising exponents, within one standard deviation.
The Hopf bifurcation is not the only way to produce a limit cycle. Another bifurcation relevant here is the infinite-period bifurcation (or SNIC bifurcation) that describes the transition from a static ordered (symmetry-breaking) state to a limit cycle [82]. To the best of our knowledge, the field-theoretical version of this bifurcation has not been studied. It has been conjectured that it does not exist in spatially extended systems of arbitrary dimension in Ref. [92] based on the instability of the static ordered state in the three-state oscillator model of Refs. [90, 91]. Our simulations suggest that a static ordered state can exist despite the presence of non-reciprocal coupling (provided that they are not fully anti-symmetric, at least in our model), but even in this case we did not find evidence of any direct transition between the static order and oscillating phases.
II.3 Stability of spatially coherent oscillations
Our work is also directly related to the question of the stability and the coherence of time-periodic phases in noisy locally-coupled classical many-body systems. By considering the nucleation and growth of droplets from fluctuations, Ref. [93, 94, 95, 96] argued that coherent oscillations are unstable in the thermodynamic limit if they arise from the spontaneous breaking of a discrete time-translation invariance (in particular when the period of oscillation is commensurate with the discrete time unit characterizing the evolution)  [^2] Note that these arguments require that the system is isotropic and that the noise is unbounded so that it can create arbitrarily large droplets [95, 96].. This is not the case in our model, where the oscillating phase arises from the spontaneous breaking of a continuous time-translation invariance (and the period is generically incommensurate). The case of spontaneously broken continuous time-translation invariance is considered in Refs. [95, 98, 96] using a mapping between the desynchronization of local oscillators and the roughening of surfaces described by the Kardar-Parisi-Zhang (KPZ) equation [99]. Based on this mapping, it is argued that temporal order can exist in $d>2$ but not in $d\leq 2$ (in isotropic systems). Note that in the mapping to the KPZ equation, the angular nature of phase variables (defined on a compact space) is neglected, which amounts to ignoring topological defects, expected not to affect the large-scale physics [98]. A compact version of KPZ equation has been studied in Refs. [100, 101, 102, 103], but not in isotropic systems in $d=3$. Similarly, Refs. [104, 87] suggested an analogy with the Mermin-Wagner theorem [105, 81] in the time domain, according to which the fluctuations of the Goldstone modes associated with broken time-translation invariance destroy order in $d\leq 2$.
References [106, 107, 108, 109, 110, 111, 112, 113] performed numerical simulations of cellular automata and coupled map lattices (see Ref. [107] for an introduction) in various space dimensions, from which it emerged that periodic oscillations are possible for $d\geq 3$ within this class of sytems [108, 114, 115]. Similar behaviors were reported in simulations of predator-prey models on lattice [116, 117, 118, 119, 120, 121, 23, 27]. These results are in agreement with the theoretical arguments reviewed above, and are are also consistent with our numerical observations. Similar questions arise in the study of the synchronization of locally coupled oscillators with randomly distributed frequencies (like in the Kuramoto model), see § IV.A of Ref. [122] and references therein, but the physics is very different because of the random frequencies.
II.4 Classical continuous time crystals
The works reviewed in the previous paragraphs are primarily concerned with the spatial coherence of the oscillations. In a nutshell, they ask whether there can be macroscopic quantities averaged over the entire system (such as the magnetization; we call these order parameters) that oscillate forever. When this is the case, we can ask another question: is the oscillation of the macroscopic order parameter temporally coherent? In other words, is it acting as a good clock? Addressing this question is another of the key motivations of our work. To illustrate the difference, let us consider a simple noisy phase oscillator
$\dot{\phi}=\omega+\sqrt{2D_{\phi}}\eta(t)$ (5)
where $\eta(t)$ is a white noise with $\langle\eta\rangle=0$ and $\langle\eta(t)\eta(t^{\prime})\rangle=\delta(t-t^{\prime})$. Here, $\phi$ can be seen as the argument of a complex number $z(t)=Ae^{i\phi(t)}$ whose amplitude $A$ is very quickly relaxing to its steady-state value. In this very simple system, we observe that the average value $\braket{\phi}$ always oscillates (averaging the equation of motion leads to $\braket{\phi(t)}=\braket{\phi(0)}+\omega t$). However, this oscillator only keeps track of time (i.e. of its initial phase) for so long. To see that, consider the correlation function $C(\tau)=\langle e^{i[\phi(t)-\phi(0)]}\rangle$ which describes the average phase that an assembly of independent clocks would have at time $t$, provided that they were synchronized at time $t=0$. Here, it behaves as 333 This can be seen by integrating $d\phi_{t}=\omega dt+\sqrt{2D_{\phi}}dW_{t}$ where $W_{t}$ is a standard Wiener process, leading to $\phi(t)-\phi(0)=\omega t+\sqrt{2D_{\phi}}W_{t}$. We find $C(\tau)=e^{i\omega t}\langle e^{i\sqrt{2D_{\phi}}W_{t}}\rangle$. As $W_{t}$ follows a normal distribution $\mathcal{N}(\mu=0,\sigma^{2}=t)$, its moment distribution function is $M_{W_{t}}(x)\equiv\mathbb{E}(e^{xW_{t}})=e^{\mu x+\sigma^{2}x^{2}/2}$ from which the result is found.
$C(\tau)\propto\cos(\omega t)e^{-t/\tau}$ (6)
where $\tau=1/D_{\phi}$, meaning that the oscillator looses track of its original phase after $\mathcal{Q}=\omega\tau$ oscillations ($\mathcal{Q}$ is the quality factor of the oscillator). In fact, any noisy oscillator behaves in this way at long times.
Thermodynamic bounds on the number $\mathcal{Q}$ of coherent oscillations have been considered [124, 125, 126, 127], and suggest that an infinite amount of energy (more precisely of entropy production) is needed to obtain infinite coherence. It is however not excluded that a macroscopic system in the thermodynamic limit can have an infinite coherence time with a finite cost per unit volume. Our numerical results suggest that this indeed happens in our model: we evaluate the coherence time of oscillations, which scales as $\tau(L)\sim L^{3}$ (in $d=3$). This points to the existence of perfectly coherent oscillations in the thermodynamic limit, in which the rate of entropy production per unit volume is finite. In this sense, our model can be seen as a classical continuous time crystal [128, 129, 130, 131, 132, 133], see also Refs. [128, 134, 135, 136, 129, 137, 138, 139] on time crystals. We conjecture that this property may extend to other systems with locally coupled oscillators that spontaneously break a continuous time-translation symmetry.
III Outline and summary
The outline of the paper is as follows: we present the non-reciprocal Ising model with general asymmetric coupling between the two species and derive its mean-field equation in Sec. IV. In Secs. V and  VIIXI we restrict ourselves to the fully anti-symmetric case in which the system exhibits $C_{4}$ symmetry. In Sec. IV we show that the mean-field equation predicts three distinct phases: disorder, time-dependent order that we dub “swap”, and static order, and analyze the bifurcations that separate the phases. Details on the Monte-Carlo simulations are give in Sec. VI. Our numerical simulations in 2D are presented in Sec. VII, suggesting that the swap phase is unstable in 2D due to spiral defects. In Sec. VIII we show that the swap phase is stable in 3D, and calculate the static critical exponents of the phase transition from disorder to swap. These critical exponents are in excellent agreement with the equilibrium XY critical exponents, which is compatible with renormalization-group studies and with the emergent continuous time-translation symmetry. The coherence time of the magnetization is shown to grow with system size suggesting a time-crystal behavior in the thermodynamic limit. In Sec. IX, we provide numerical evidence and theoretical arguments that suggest that the static-order phase is destroyed in any finite dimension due to droplet growth. In Sec. X we discuss how the system coarsens into the swap and static-order phases. In Sec. XII we allow the presence of inter-species reciprocal couplings, which reduces the symmetry to $\mathbb{Z}_{2}$, and accounts for a general asymmetric interaction. We present mean-field analysis and 2D and 3D simulation results. These results reveal that the swap phase stability is unchanged with respect to the fully asymmetric case (i.e. it is stable in 3D but not in 2D). However, static order can be retrieved by an intriguing droplet-capture mechanism. Contrary to the mean-field prediction, we find no evidence of a direct transition between the swap and the static-order phase. We discuss our results and draw conclusions in Sec. XIII.
IV The model
In the equilibrium Ising model, adjacent on-lattice spins ($\sigma=\pm 1$) stochastically align with one another through an energy function. Its dynamics is given by kinetic Ising models which, while lacking uniqueness, always operate by minimizing the global free energy. In what follows, we generalize these kinetic Ising models to the case of non-reciprocal interactions where such global potential function does not exist. To this aim, we define a Glauber dynamics [140] where each spin minimizes its own selfish energy. Similar formulations were used among the physical sciences [65, 52, 35, 141, 142, 143, 20, 144, 145, 146, 6].
IV.1 The non-reciprocal (NR) Ising model
We consider two spin species, labeled by Greek indices $\alpha=A,B$, located on a $d$-dimensional cubic lattice of linear size $L$. Each site $i\in\{1,...,L^{d}\}$ contains one A-spin and one B-spin (Fig. 1, left).
The state of the system is described by a vector
$\vec{\sigma}=\{\sigma^{A}_{1},...,\sigma^{A}_{L^{d}},\sigma^{B}_{1},...,\sigma^{B}_{L^{d}}\}$ (7)
in which each spin $\sigma^{\alpha}_{i}=\pm 1$.
We consider transition rates given by the generalized Glauber dynamics [140]
$w(F^{\alpha}_{i}\vec{\sigma}|\vec{\sigma})=\frac{1}{2\tau}\left[1-\tanh\left(\frac{\Delta E^{\alpha}_{i}}{2k_{B}T}\right)\right]$ (8)
where $\tau$ is the single spin-flip time-scale, $T$ is the temperature, $k_{B}$ the Boltzmann constant and
$\Delta E^{\alpha}_{i}=E^{\alpha}_{i}(F^{\alpha}_{i}\vec{\sigma})-E^{\alpha}_{i}(\vec{\sigma})$ (9)
is the difference in selfish energy between the states. Here $F^{\alpha}_{i}$ is a single-spin-flipping operator, i.e., when acted on $\vec{\sigma}$ it flips the $\alpha$-species spin on site $i$ from $\sigma^{\alpha}_{i}$ to $-\sigma^{\alpha}_{i}$ while keeping all the other spins unchanged. For instance,
$\begin{split}&F_{i}^{A}\{\sigma_{1}^{A},...,\sigma_{i}^{A},...,\sigma_{L^{d}}^{A},\sigma_{1}^{B},...,\sigma_{L^{d}}^{B}\}\\ &\quad=\{\sigma_{1}^{A},...,-\sigma_{i}^{A}...,\sigma_{L^{d}}^{A},\sigma_{1}^{B},...,\sigma_{L^{d}}^{B}\}\end{split}$ (10)
Finally, $w(\vec{\sigma}|\vec{\sigma}^{\prime})$ is the transition rate from configuration $\vec{\sigma}$ to $\vec{\sigma}^{\prime}$.
The selfish energy is
$E^{\alpha}_{i}=-J\sum_{j\,{\rm nn\,of}\,i}\sigma_{i}^{\alpha}\sigma_{j}^{\alpha}-K_{\alpha\beta}\sigma_{i}^{\alpha}\sigma_{i}^{\beta},$ (11)
where
$K_{\alpha\beta}=\left(\begin{array}[]{cc}0&K_{+}+K_{-}\\ K_{+}-K_{-}&0\end{array}\right)$ (12)
and $J,K_{+},K_{-}>0$. The sum in Eq. (11) runs over nearest neighbors of $i$ and summation over spin indices $\beta$ is implied. The selfish energy indicates that spins of the same species tend to align with their neighbors, while spins of different species interact non-reciprocally (see Fig. 1, right): spins $A$ tend to align with spins $B$, whereas spins $B$ either tend to align with spins $B$ but to a lesser extent (if $K_{+}>K_{-}$), or tend to anti-align with spins $A$ (if $K_{+}<K_{-}$). The latter case will be the main focus of this paper as it leads to a rich phenomenology related to limit cycles. For simplicity, note that inter-species interactions only occur on the same site.
Finally, in the continuous-time limit, the probability $P(\vec{\sigma},t)$ to observe a specific configuration of spins $\vec{\sigma}$ at time $t$ evolves according to the master equation
$\partial_{t}P(\vec{\sigma},t)$ $=-\sum_{i,\alpha}w(F^{\alpha}_{i}\vec{\sigma}|\vec{\sigma})P(\vec{\sigma},t)$ (13)
$+\sum_{i,\alpha}w(\vec{\sigma}|F^{\alpha}_{i}\vec{\sigma})P(F^{\alpha}_{i}\vec{\sigma},t).$
[…]
Fig. 1: Schematic drawing of the non-reciprocal (NR) Ising model. The model includes two species per site with two interaction types: (i) reciprocal nearest-neighbors interaction of strength $J$ between same-species spins and (ii) non-reciprocal same-site interaction of strength $K_{+}+K_{-}$ for species A (blue) and $K_{+}-K_{-}$ for species B (red), between spins of different species.
Fig. 2: Schematic drawing of the mapping between the state of two spins located on the same site ($\sigma^{A}$ and $\sigma^{B}$) into its corresponding angle variable $\theta$. The color associated to each quadrant is used throughout the paper to indicate the corresponding spin state. The direction of the cycle is determined by the sign of $K_{-}$ (throughout the paper $K_{-}>0$ which implies clockwise direction).
IV.2 Broken detailed balance
A Markov chain with rates $w$ satisfies detailed balance if it has a stationary distribution $P_{\text{s}}$ such that
$\frac{w(\vec{\sigma}\to\vec{\sigma}^{\prime})}{w(\vec{\sigma}^{\prime}\to\vec{\sigma})}=\frac{P_{\text{s}}(\vec{\sigma}^{\prime})}{P_{\text{s}}(\vec{\sigma})}.$ (14)
Let us assume that our kinetic model possesses an (unknown) stationary distribution $P_{\text{s}}(\vec{\sigma})$. Considering a loop $\vec{\sigma}_{1}\to\vec{\sigma}_{2}\to\cdots\to\vec{\sigma}_{n}$ in state space, as well as the same loop travelled in opposite direction, detailed balance implies [147]
$\frac{w(\vec{\sigma}_{1}\to\vec{\sigma}_{2})}{w(\vec{\sigma}_{1}\to\vec{\sigma}_{n})}\cdot\frac{w(\vec{\sigma}_{2}\to\vec{\sigma}_{3})}{w(\vec{\sigma}_{n}\to\vec{\sigma}_{n-1})}\cdots\frac{w(\vec{\sigma}_{n}\to\vec{\sigma}_{1})}{w(\vec{\sigma}_{2}\to\vec{\sigma}_{1})}=1.$ (15)
[…]
$\begin{split}w\left(\uparrow\uparrow\to\uparrow\downarrow\right)&=f(\tilde{J}+\tilde{K}_{+}-\tilde{K}_{-})\\ w\left(\uparrow\downarrow\to\downarrow\downarrow\right)&=f(\tilde{J}-\tilde{K}_{+}-\tilde{K}_{-})\\ w\left(\downarrow\downarrow\to\downarrow\uparrow\right)&=f(-\tilde{J}+\tilde{K}_{+}-\tilde{K}_{-})\\ w\left(\downarrow\uparrow\to\uparrow\uparrow\right)&=f(-\tilde{J}-\tilde{K}_{+}-\tilde{K}_{-})\\ w\left(\uparrow\uparrow\to\downarrow\uparrow\right)&=f(\tilde{J}+\tilde{K}_{+}+\tilde{K}_{-})\\ w\left(\downarrow\uparrow\to\downarrow\downarrow\right)&=f(\tilde{J}-\tilde{K}_{+}+\tilde{K}_{-})\\ w\left(\downarrow\downarrow\to\uparrow\downarrow\right)&=f(-\tilde{J}+\tilde{K}_{+}+\tilde{K}_{-})\\ w\left(\uparrow\downarrow\to\uparrow\uparrow\right)&=f(-\tilde{J}-\tilde{K}_{+}+\tilde{K}_{-})\end{split}$ (16)
where $f(x)\equiv\left(1-\tanh(x)\right)/(2\tau)$ and we defined the normalized coupling coefficients $\tilde{K}_{-}=K_{-}/(k_{B}T)$, $\tilde{K}_{+}=K_{+}/(k_{B}T)$ and $\tilde{J}\equiv 2dJ/(k_{B}T)$. Using the identity $(1+\tanh(x))/(1-\tanh(x))={\rm e}^{2x}$ we find that the forward and reverse cycles satisfy
$\!\!\!\frac{w\left(\uparrow\uparrow\to\uparrow\downarrow\right)w\left(\uparrow\downarrow\to\downarrow\downarrow\right)w\left(\downarrow\downarrow\to\downarrow\uparrow\right)w\left(\downarrow\uparrow\to\uparrow\uparrow\right)}{w\left(\uparrow\downarrow\to\uparrow\uparrow\right)w\left(\downarrow\downarrow\to\uparrow\downarrow\right)w\left(\downarrow\uparrow\to\downarrow\downarrow\right)w\left(\uparrow\uparrow\to\downarrow\uparrow\right)}=e^{8\tilde{K}_{-}}$ (17)
which means that detailed balance is broken in our model whenever $\tilde{K}_{-}\neq 0$.
IV.3 The mean-field (MF) equation
The average magnetization of a single spin at time $t$ is given by
$\langle\sigma^{\alpha}_{i}\rangle\left(t\right)=\sum_{\vec{\sigma}}\sigma^{\alpha}_{i}P({\vec{\sigma}},t).$ (18)
Using Eq. (13) and changing summation variables we get
$\frac{\rm{d}}{{\rm d}t}\langle\sigma^{\alpha}_{i}\rangle=-2\langle\sigma^{\alpha}_{i}w(F^{\alpha}_{i}\vec{\sigma}|\vec{\sigma})\rangle.$ (19)
Using Eq. (8) and substituting the selfish energy (Eq. (11)) we obtain
$\begin{split}&\tau\frac{{\rm d}}{{\rm d}t}\langle\sigma_{i}^{\alpha}\rangle=-\langle\sigma_{i}^{\alpha}\rangle\\ &+\left\langle\tanh\left[\frac{J}{k_{B}T}\sum_{j\,{\rm nn\,of}\,i}\sigma_{j}^{\alpha}+\frac{K_{\alpha\beta}}{k_{B}T}\sigma_{i}^{\beta}\right]\right\rangle\end{split}$ (20)
where we have further used the identity $x\tan(xb)=\tan(b)$ for $x=\pm 1$. Applying the mean-field (MF) approximation, $\langle g(\vec{\sigma})\rangle=g(\langle\vec{\sigma}\rangle)$ and defining the average magnetization as $\langle\sigma^{\alpha}_{i}\rangle=m^{\alpha}_{i}$ yields
$\partial_{t}m_{i}^{\alpha}=-\frac{m_{i}^{\alpha}}{\tau}+\frac{1}{\tau}\tanh\left(\frac{J\sum_{j\,{\rm nn\,of}\,i}m_{j}^{\alpha}+K_{\alpha\beta}m_{i}^{\beta}}{k_{B}T}\right).$ (21)
We now assume that the average magnetization changes over length scales much larger than the lattice spacing and write it as a continuum variable $m_{i}^{\alpha}=m_{\alpha}({\vec{r}},t)$. We thus get
$\partial_{t}m_{\alpha}=-\frac{m_{\alpha}}{\tau}+\frac{1}{\tau}\tanh\left(\frac{2dJm_{\alpha}+K_{\alpha\beta}m_{\beta}+Ja^{2}\nabla^{2}m_{\alpha}}{k_{B}T}\right)\;,$ (22)
where $a$ is the lattice spacing. Finally, upon rescaling time by $\tau$ and space by $a$ we obtain
$\partial_{t}m_{\alpha}=-m_{\alpha}+\tanh\left[\tilde{J}m_{\alpha}+\tilde{K}_{\alpha\beta}m_{\beta}+D\nabla^{2}m_{\alpha}\right]$ (23)
where $D\equiv J/(k_{B}T)$ and $\tilde{K}_{\alpha\beta}\equiv K_{\alpha\beta}/(k_{B}T)$. Equation (23) is the MF equation of the general NR Ising model. Note that it is invariant under the $\mathbb{Z}_{2}$ symmetry $(m_{A},m_{B})\mapsto(-m_{A},-m_{B})$.
V The fully anti-symmetric case
In this section as well as in Secs. VIIXI, we will study the fully anti-symmetric case, in which $K_{+}=0$. To simplify the notations, we write $K_{-}\equiv K$ (and $\tilde{K}_{-}\equiv\tilde{K}$). In this case, the MF equation reads
$\partial_{t}m_{\alpha}=-m_{\alpha}+\tanh\left[\tilde{J}m_{\alpha}+\tilde{K}\varepsilon_{\alpha\beta}m_{\beta}+D\nabla^{2}m_{\alpha}\right]$ (24)
where $\varepsilon_{\alpha\beta}$ is the Levi-Civita symbol. Equation (24) has the additional symmetry $(m_{A},m_{B})\mapsto(-m_{B},m_{A})$, so overall it is invariant under $C_{4}$.
V.1 Connection with locally-coupled oscillators
In the fully anti-symmetric case, the dynamics of $\theta_{i}$ is described by the forward and backward moves,
$\theta_{i}\to\theta_{i}+\frac{\pi}{2}\,\,\,\,{\rm with\,\,rate}\,\,\,\,w^{+}$ (25a)
$\theta_{i}\to\theta_{i}-\frac{\pi}{2}\,\,\,\,{\rm with\,\,rate}\,\,\,\,w^{-}$ (25b)
where the rates are
$w^{\pm}=\frac{1}{2\tau}\left[1-\tanh\left(\frac{\Delta E(\vec{\theta})}{2K_{B}T}\pm\tilde{K}\right)\right]$ (26)
and
$E(\vec{\theta})=-2J\sum_{\langle i,j\rangle}\cos\left(\theta_{i}-\theta_{j}\right).$ (27)
Equations (25)-(27) can be viewed as describing locally-coupled oscillators with internal frequency and discrete states. In this sense, they share similarities with the Wood et al. cyclic model [91, 90, 92] albeit in that model oscillators are restricted to forward moves (backward moves are prohibited). For $K=0$, our model reaches an equilibrium state that is equivalent to the clock model (vector Potts model) for $q=4$ [148, 149].
Fig. 3: Mean-field (MF) flow in $\left(m_{A},m_{B}\right)$ space. Arrows correspond to the vector $\vec{v}=(\partial_{t}m_{A},\partial_{t}m_{B})$ calculated from the expanded MF equation (top row, Eq. (30)) and full MF equation (bottom row, Eq. (24) with $D=0$). The limit cycle in the swap phase is highlighted by a semi-transparent blue line. Non-reciprocal inter-species coupling is $\tilde{K}=0.4$, and reciprocal coupling between same-species nearest neighbours is (a,d) $\tilde{J}=0.8$ , (b,e), $\tilde{J}=1.3$, and (c,f) $\tilde{J}=1.8$, showing the disorder, swap, and static-order phase, respectively. While the exact flow varies between expanded (top) and full (bottom) MF equations, the qualitative behavior is the same.
V.2 Analysis of the bifurcations
In the ${m_{\alpha}\ll 1}$ limit and for small gradients, the rescaled MF equation in the fully anti-symmetric case (Eq. (24)) can be approximated by
$\partial_{t}m_{\alpha}=$ $-(1-\tilde{J})m_{\alpha}+\tilde{K}\varepsilon_{\alpha\beta}m_{\beta}+D\nabla^{2}m_{\alpha}$ (28)
$-\frac{1}{3}\left(\tilde{J}m_{\alpha}+\tilde{K}\varepsilon_{\alpha\beta}m_{\beta}\right)^{3}.$
For $\tilde{K}=0$, Eq. (28) reduces to two uncoupled MF equations associated with $\phi^{4}$ Ginzburg-Landau theory. For any $\tilde{K}\neq 0$ the equations can not be derived from a Hamiltonian dynamics nor a gradient descent, which corresponds to its far-from-equilibrium nature. Note that due to the lack of rotational symmetry in the cubic term, Eq. (28) also differs from the Complex Ginzburg-Landau (CGL) equation [84, 150] given by
$\partial_{t}\psi=a\psi+b\nabla^{2}\psi+c|\psi|^{2}\psi$ (29)
where $\psi({\vec{r}},t)$ is a complex field and $a$, $b$ and $c$ are complex numbers. However, Eq. (28) becomes CGL-like in special limits as discussed in Sec. V.2.
A linear stability analysis shows that any homogeneous solution, $m_{\alpha}({\vec{r}},t)\equiv m_{\alpha}(t)$, that constitutes a stable fixed point of the spatially homogeneous case ($D=0$) remains stable in the inhomogeneous case ($D>0$). Therefore, in what follows we analyze the spatially homogeneous MF equation given by
$\partial_{t}m_{\alpha}=$ $-(1-\tilde{J})m_{\alpha}+\tilde{K}\varepsilon_{\alpha\beta}m_{\beta}$ (30)
$-\frac{1}{3}\left(\tilde{J}m_{\alpha}+\tilde{K}\varepsilon_{ab}m_{\beta}\right)^{3}.$
Figure 3 shows the phase-space velocity field $\vec{v}=(\partial_{t}m_{A},\partial_{t}m_{B})$ from Eq. (30) along with the fixed points which are obtained by numerically solving the $\partial_{t}m_{A}=\partial_{t}m_{B}=0$ equations. The stability of the fixed points (shown in the figure as orange, pink, and brown for stable, saddle, and unstable, respectively) is determined from the eigenvalues of the Jacobian matrix $\mathcal{J}$ of the linearized equation $\partial_{t}\delta\vec{m}=\mathcal{J}\delta\vec{m}$ where linearization is performed around the fixed point.
Three distinct phases are found: a disordered phase in which ${m_{A}=m_{B}=0}$ is the only stable fixed point (Fig. 3a), a limit-cycle phase that we dub “swap phase” as both species repeatedly flip their magnetizations (Fig. 3b), and a static-order (ferromagnetic) phase in which, depending on initial conditions, the system ends up in one out of four stable fixed points with non-zero $m_{A}$ and $m_{B}$ (Fig. 3c).
The MF phase diagram is shown in Fig 4, as a function of the couplings $\tilde{J}$ and $\tilde{K}$. The phases are determined by the fixed points and their stability as described above, but for generality sake we used the full non-linear homogeneous equation, namely Eq. (24) with $D=0$. When non-reciprocal interactions are turned off (${\tilde{K}=0}$), a pitchfork bifurcation at ${\tilde{J}_{c}\equiv 1}$ (red point in Fig. 4) separates the disordered phase (in blue) from the static-order phase (in green), as in the usual Ising model. When non-reciprocal interactions are present (${\tilde{K}\neq 0}$), the swap phase arises (in pink) and is separated from the disordered state by a Hopf bifurcation at $\tilde{J_{c}}=1$ (green line), and from the static-order phase by another bifurcation known as a saddle-node on an invariant circle (SNIC) bifurcation [82, 64] (blue line ${\tilde{K}_{\rm SNIC}(\tilde{J})}$). We analyze each one of these bifurcations below.
Fig. 4: Mean-field phase diagram. The three phases shown on the diagram are (i) disorder, (ii) swap and (iii) static order. The lines separating the phases are Hopf bifurcation (thin green) and SNIC bifurcation (thick blue). Red dot is the Pitchfork bifurcation. Obtained from numerical analysis of Eq. (24) with $D=0$.
V.2.1 Pitchfork bifurcation
For ${\tilde{K}=0}$, Eq. (30) leads to the standard pitchfork bifurcation at ${\tilde{J}_{\rm PF}=1}$. The ${m_{A}=m_{B}=0}$ point becomes unstable and four stable fixed points
$m_{A}=\pm m_{B}=\pm[3(\tilde{J}-1)/\tilde{J}^{3}]^{1/2}$ (31)
emerge, corresponding to the ordered phase.
V.2.2 Hopf bifurcation
For $\tilde{K}\neq 0$, the eigenvalues of $\mathcal{J}$ at $m_{A}=m_{B}=0$ are complex,
$\lambda_{1,2}=-1+\tilde{J}\pm i\tilde{K}.$ (32)
This means that there is a Hopf bifurcation at ${\tilde{J}_{\rm Hopf}=1}$: the ${m_{A}=m_{B}=0}$ point become unstable and a stable limit cycle emerges, corresponding to the swap phase.
The limit cycle is better understood in polar coordinates, ${m_{A}=r\cos\theta}$ and ${m_{B}=r\sin\theta}$ where Eq. (30) becomes
$\partial_{t}r$ $=-\left(1-\tilde{J}\right)r-\frac{1}{4}\tilde{J}^{3}\bigg{[}1+\kappa^{2}$
$+\left(\frac{1}{3}-\kappa^{2}\right)\cos(4\theta)+\left(\kappa-\frac{1}{3}\kappa^{3}\right)\sin(4\theta)\bigg{]}r^{3}$ (33)
$\partial_{t}\theta$ $=-\tilde{K}+\frac{1}{4}\tilde{J}^{3}\bigg{[}\kappa+\kappa^{3}$
$+\left(\frac{\kappa^{3}}{3}-\kappa\right)\cos(4\theta)+\left(\frac{1}{3}-\kappa^{2}\right)\sin(4\theta)\bigg{]}r^{2}\;,$ (34)
with $\kappa\equiv\tilde{K}/\tilde{J}$.
The limit cycle leads to anharmonic oscillations, where both $r$ and $\dot{\theta}$ depend on $\theta$. Its shape has a rotational symmetry of order 4 due to the $C_{4}$ symmetry of the mean-field equation. However, very close to the bifurcation, i.e. $\tilde{J}-1\ll 1$, the oscillations are harmonic [151]. This can be seen as follows.
The equation for $\theta$ in the ${\tilde{J}-1\ll 1}$ limit is
$\partial_{t}\theta=-\tilde{K}+\mathcal{O}(\tilde{J}-1)\;,$ (35)
where we used the fact that ${r^{2}\sim 1-\tilde{J}}$. The angle $\theta$ then has the solution ${\theta=\theta_{0}+\tilde{K}t}$, and the equation for $r$ becomes
$\partial_{t}r$ $=-\left(1-\tilde{J}\right)r-\frac{1}{4}\bigg{[}1+\tilde{K}^{2}+\left(\frac{1}{3}-\tilde{K}^{2}\right)\cos(4\tilde{K}t)$ (36)
$-\left(\tilde{K}-\frac{1}{3}\tilde{K}^{3}\right)\sin(4\tilde{K}t)\bigg{]}r^{3}+\mathcal{O}\left((\tilde{J}-1)^{5/2}\right)$
where we have set $\theta_{0}$ to zero. Equation (36) can be solved analytically. In steady state ($t\to\infty$ limit), its solution reads
$r\left(t\right)=2\sqrt{\frac{\tilde{J}-1}{1+\tilde{K}^{2}}}\times\frac{1}{\sqrt{1-\left(\tilde{J}-1\right)f\left(\tilde{J},\tilde{K},t\right)}}$ (37)
with
$f\left(\tilde{J},\tilde{K},t\right)$ $=\frac{1-\tilde{J}+3\left(\tilde{J}-3\right)\tilde{K}^{2}+2\tilde{K}^{4}}{3\left(1+\tilde{K}^{2}\right)\left[\left(\tilde{J}-1\right)^{2}+4\tilde{K}^{2}\right]}\cos\left(4\tilde{K}t\right)$ (38)
$-\frac{\tilde{K}\left(5-3\tilde{J}+\left(\tilde{J}-7\right)\tilde{K}^{2}\right)}{3\left(1+\tilde{K}^{2}\right)\left[\left(\tilde{J}-1\right)^{2}+4\tilde{K}^{2}\right]}\sin\left(4\tilde{K}t\right)$
Taking the limit $\tilde{J}-1\ll 1$ once more $(\tilde{J}-1)f(\tilde{J},\tilde{K},t)$ goes to zero and we arrive at
$r\left(t\right)=2\sqrt{\frac{\tilde{J}-1}{1+\tilde{K}^{2}}}$ (39)
which is a circular motion of constant angular velocity in phase space. This harmonic solution can be rationalized by looking at the two times scales corresponding respectively to the angular and radial motion close to the bifurcation (described by Eqs. (35) and (36), respectively). The time scale for the angular motion is $T_{\theta}\sim\tilde{K}^{-1}$ while the time scale of the radial motion is determined by the relaxation time of $r(t)$ to its new minimum when $\theta$ changes. Thus, it is found by linearizing the equation for $r$ around 00: this yields $T_{r}\sim(\tilde{J}-1)^{-1}$. The two time scales $T_{\theta}$ and $T_{r}$ can be viewed as “oscillation time-scale” and “alignment time-scale”, respectively. Since $T_{r}\gg T_{\theta}$, the angular motion is much faster and there is not enough time for $r$ to reach a new minimum when $\theta$ changes. Thus, only the $\theta$-averaged value of the right hand side of Eq. (V.2.2) matters. This is supported by Fig. 5 which shows typical phase-space trajectories in the swap phase: as $\tilde{J}-1$ increases, the motion indeed loses its rotational symmetry and become more “square-like”.
Fig. 5: Swap phase at different distances from the Hopf bifurcation. Numerical solutions of the spatially homogeneous MF dynamics, Eq. (30). (a-c) $m_{B}$ vs. $m_{A}$. (d-f) $m_{A}$ vs. $t$. From left to right: $\tilde{J}=1.0002$, $1.001$, and $1.0028$. In all plots $\tilde{K}=0.001$. Red dot is the initial state and dashed black line is the steady state predicted from Eq. (39), valid in the $\tilde{J}-1\ll 1$ limit.
The above result is a consequence of a more general theorem. A reaction-diffusion equation that undergoes a Hopf bifurcation such as the MF evolution Eq. (24) can be reduced, close to the bifurcation, to the CGL equation (i.e., the CGL equation is its normal form). The transformation to the CGL equation guarantees that the steady-state solution of the non-transformed set of equations is an elliptical limit cycle (or in special cases like here, circular) [151]. Because of this reduction, we expect the disorder-to-swap phase transition in our model to exhibit the same critical exponents as the Hopf bifurcation in the CGL equation, which has been studied in Refs. [85, 86].
V.2.3 Infinite-period (SNIC) bifurcation
Let us now start from the limit-cycle swap phase. Upon increasing $\tilde{J}$ or decreasing $\tilde{K}$, the system goes through a ($C_{4}$ equivariant) saddle-node on an invariant circle (SNIC) bifurcation, where four pairs of stable and saddle fixed points emerge from the limit cycle (pink and orange points in Fig. 3c,e). This bifurcation is also known under several other names: infinite-period bifurcation (because the period of oscillation diverges at the bifurcation point), saddle-node on a limit cycle (SNLC), saddle-node infinite-period (SNIPER), etc. Once the stable fixed points emerge, the system is in a static-order phase.
Noticing that the infinite-period bifurcation occurs when $\tilde{K}$ and $\tilde{J}-1$ are comparable (see Fig. 4), we now take the limit $\epsilon\to 0$ with $\tilde{J}-1\sim\epsilon$ and $\tilde{K}\sim\epsilon$. At first order in $\epsilon$, Eqs. (V.2.2) and (34) become
$\partial_{t}r$ $=\left(\tilde{J}-1\right)r-\frac{1}{4}\left(1+\frac{1}{3}\cos(4\theta)\right)r^{3}\;,$ (40a)
$\partial_{t}\theta$ $=-\tilde{K}+\frac{1}{12}\sin(4\theta)r^{2}.$ (40b)
In a static-order state, $\partial_{t}r=\partial_{t}\theta=0$, with $r$ given by
$r=2\sqrt{\frac{\tilde{J}-1}{1+\frac{1}{3}\cos(4\theta)}}\;.$ (41)
The equation for $\theta$ then becomes
$\partial_{t}\theta=-\tilde{K}+\frac{\sin(4\theta)}{3+\cos(4\theta)}\left(\tilde{J}-1\right).$ (42)
Since $\sin(4\theta)/(3+\cos(4\theta))$ is bounded between $-1/(2\sqrt{2})$ and $1/(2\sqrt{2})$, a stationary state can only exist if ${\tilde{K}>(\tilde{J}-1)/(2\sqrt{2})}$. Therefore, in the regime $\tilde{J}-1\ll 1$, the SNIC bifurcation line corresponds to
$\tilde{K}_{\rm SNIC}=\frac{\tilde{J}_{\rm SNIC}-1}{2\sqrt{2}}.$ (43)
Note that, in Fig. 4, the thick blue line corresponding to the SNIC bifurcation was obtained numerically for the full nonlinear MF dynamics Eq. (24). Thus, our Eq. (43) only predicts the tangent of this thick blue line close to the bifurcation point $\tilde{J}_{\rm PF}=1$.
Finally, we find the stable points emerging at the SNIC bifurcation by extremizing $\sin(4\theta)/(3+\cos(4\theta))$ with respect to $\theta$. Those are
$\theta_{n}=\left[-\tan^{-1}\left(2\sqrt{2}\right)+\pi\left(2n+1\right)\right]/4$ (44)
for $n=0,1,2,3$, and their corresponding $r$ can be obtained from Eq. (41) as
$\left(m_{A},m_{B}\right)=3\sqrt{\frac{\tilde{J}-1}{2}}\left(\cos(\theta_{n}),\sin(\theta_{n})\right).$ (45)
Approaching the SNIC bifurcation from the side of the limit cycle (swap phase), $\partial_{t}\theta$ approaches zero when the system passes at $\theta=\theta_{n}$. As a result, the oscillation’s period diverges at the bifurcation (see Fig. 5d-f, where the period increases as we approach the bifurcation). This stands in contrast with the Hopf bifurcation in which the period of oscillations remains finite.
VI Monte-Carlo simulations
To go beyond the MF description, we perform large-scale Monte-Carlo simulations of the NR Ising model in both two and three dimensions (results in one dimension are summarized in Appendix D).
In our 2D discrete-time Monte-Carlo (MC) simulations, we use the standard Glauber algorithm: at each MC step a spin $\sigma_{i}^{\alpha}$ is chosen at random and then flipped with probability
$p_{i}^{\alpha}=\tau w(F^{\alpha}_{i}\vec{\sigma}|\vec{\sigma})$ (46)
(with $w$ defined in Eq. (8)).
Since three-dimensional simulations are much more costly, in 3D we update spins using a checkerboard algorithm to implement Eq. (46) instead of random picking. In the checkerboard algorithm non-interacting spins are flipped in parallel [152]. To generalize the checkerboard approach to the two-species case, we defined the quantity $\mathcal{P}=x_{1}+x_{2}+x_{3}+\alpha$ where $x_{n}$ is the integer position in the $n$-th dimension of the cubic lattice, and $\alpha=0,1$ labels the two species. Spins with the same parity of $\mathcal{P}$ (even or odd) are then updated simultaneously. This allows us both to parallelize the simulation and avoid random number generation in choosing the spins. We did not observe qualitative changes when changing the update rules from random picking to paralleled checkerboard, as well as when using other update rules, see Appendix A.
Unless mentioned otherwise, we initialize the system in an ordered state where all spins of the same species are either up or down, and let the system evolve until a steady-state is reached. Time is measured in simulation sweeps, where a single sweep is defined as $2\times L^{d}$ (the number of spins) MC steps.
Transitions from disorder to time-dependent order and from time-dependent order to stationary order can be quantified by two order parameters. First, the synchronization order parameter
$R\equiv\left\langle s\right\rangle_{t,\Omega}\,\,\,\,\text{with}\,\,\,s\equiv\sqrt{\frac{M_{A}^{2}+M_{B}^{2}}{2}}$ (47)
quantifies how much spins are aligned with each other within each species. Here, $M_{\alpha}=\sum_{i}\sigma_{i}^{\alpha}/L^{d}$ are the total magnetizations and the average $\langle...\rangle_{t,\Omega}$ is over time ($t$) and realizations ($\Omega$). Note that $s$ can also be written as
$s=\frac{1}{L^{d}}\bigg{|}\sum_{j}{\rm e}^{i\theta_{j}}\bigg{|},$ (48)
similarly to the order parameter quantifying the synchronization of oscillations in, for example, the Kuramoto model [122]. Second, the phase-space angular momentum
$\mathcal{L}\equiv\langle\ell\rangle_{t,\Omega}\,\,\,{\rm with}\,\,\,\ell$ $\equiv M_{B}\partial_{t}M_{A}-M_{A}\partial_{t}M_{B}$
$=M_{B,t}M_{A,t+1}-M_{A,t}M_{B,t+1}$ (49)
quantifies the time-dependence of the magnetizations $M_{A}$ and $M_{B}$. It is related to the entropy production [153] close to the Hopf bifurcation. In mean-field, the synchronization order parameter defined in Eq. (47) is
$R_{\rm MF}=\frac{r}{\sqrt{2}}$ (50)
and the phase-space angular momentum defined in Eq. (VI) is
$\mathcal{L}_{\rm MF}=-r^{2}\dot{\theta}.$ (51)
The sign is chosen so that $\mathcal{L}_{\rm MF}$ is positive when $K$ is positive. In Fig. 6, $R_{\rm MF}$ and $\mathcal{L}_{\rm MF}$, calculated from a numerical stationary solution of Eq. (30), are plotted as a function of $\tilde{J}$ for a fixed value of $\tilde{K}$. It is shown that $R_{\rm MF}$ goes from zero to non-zero values when crossing the Hopf bifurcation at $\tilde{J}=1$. When crossing the SNIC bifurcation, $R_{\rm MF}$ remains finite, but there is a sharp change in the slope: the slope increases when approaching the bifurcation and decreases in the static order regime. Conversely, $\mathcal{L}_{\rm MF}$ is identically zero in both the disordered and static-order phases, but deviates from zero in the swap phase. We now analyze how these MF results hold in finite-dimensional systems using the Monte-Carlo simulations.
Fig. 6: Mean-field order parameters. (a) The synchronization order parameter $R_{\rm MF}$ (Eq. (50)) and (b) the phase-space angular momentum $\mathcal{L}_{\rm MF}$ (Eq. (51)), shown as a function of $\tilde{J}$ and for fixed $\tilde{K}=0.1$. The variables $r(t)$ and $\theta(t)$ are calculated from numerical solutions of Eq. (30) in steady state.
Fig. 7: Swap phase in finite 2D systems. The magnetizations $M_{A}$ (blue) and $M_{B}$ (red) as a function of time, for increasing system sizes (from top to bottom: $L=20,50,150,200$). Obtained from MC simulations with the parameters: $\tilde{J}=2.8$, $\tilde{K}=0.3$.
VII Destruction of the swap phase in 2D by spiral defects
[…]
In small 2D systems, simulations show states in which $M_{A}$ and $M_{B}$ oscillate in time (like in the MF swap phase). These oscillations are irregular, and typically occur through nucleation of droplets of opposite magnetization by the “unsatisfied” species, in alternating order (see Movie 1).
[…]
Despite having 4-fold rather than full rotational symmetric shape, the spiral waves observed in the NR Ising model appear to behave similarly to those emerging in the CGL equation [84, 156, 157, 98, 101, 102, 158, 159, 160, 161, 162]. Defects are created and annihilated in pairs and opposite topological charges ($+$ and -- defects) lead to opposite angular velocities (see Movie 1). Finally we note that in mean-field, a homogeneously oscillating system is stable, but given sufficiently strong added noise spirals emerge, as in our MC simulations (see Appendix C).
Quantitatively, the absence of a swap phase in the thermodynamic limit can be seen from the behavior of the order parameters $R$ and $\mathcal{L}$ (Eq. (47) and (VI), respectively) as system size increases. In Fig. 9 color maps of $R$ and $\mathcal{L}$ are shown as a function of $\tilde{J}$ and $\tilde{K}$ for both $L=20$ and $L=80$. For $L=20$, there is a region in which both $R$ and $\mathcal{L}$ are non-zero, compatible with the swap phase. However, for $L=80$, both $R$ and $\mathcal{L}$ diminish in this region, indicating instead a disorder phase. Figure 10 shows the order parameters as function of $\tilde{J}$ for fixed $\tilde{K}$, at different system sizes. The intermediate region ($1.5\lesssim\tilde{J}\lesssim 3$) corresponds to the finite-system swap state, but both $R$ and $\mathcal{L}$ approach zero as $L$ increases. In addition, the putative critical $\tilde{J}$ marking the transition from disorder to swap depends on the system size, signifying the absence of a well-defined phase transition in the thermodynamic limit.
To conclude, our numerical simulations indicate that the MF swap phase is destroyed by spiral defects in 2D where it is replaced by a disordered state.
Fig. 9: Instability of the swap phase in 2D shown by the order parameters. Color map of the (a) synchronization order parameter $R$ and (b) phase-space angular momentum $\mathcal{L}$, as a function of $\tilde{J}$ and $\tilde{K}$, for a 2D system with $L=20$. (c-d) The same with $L=80$. Obtained from MC simulations.
Fig. 10: No phase transition from disorder to swap in 2D. (a) $R$ and (b) $\mathcal{L}$, shown as a function of $\tilde{J}$ for $\tilde{K}=0.3$, for different linear system size $L$. Dashed red line shows upper bound for $R$ and $\mathcal{L}$ for points that did not converge during simulation running time. Both $R$ and $\mathcal{L}$ approach zero as $L$ increases.
VIII Existence of a stable swap phase in 3D
Fig. 11: Time evolution of a stable swap phase in 3D. (a) Kymograph of $\theta$ for a single 1D row in the 3D system. (b) $M_{A}$, $M_{B}$ and $s=\sqrt{\left(M_{A}^{2}+M_{B}^{2}\right)/2}$ as a function of time. Obtained from MC simulation with parameters: $\tilde{J}=1.5$, $\tilde{K}=0.1$ and $L=160$.
In 3D, numerical simulations suggest that a stable swap phase does exist. Its behavior is demonstrated in Fig. 11 where we show a kymograph of the $\theta$-field of a single 1D row in the 3D lattice as well as plot the total magnetizations and $s$ (Eq. (47)) as a function of time. A full visualization of this phase is shown in Movie 2. Unlike the 2D droplet-induced swap persisting at finite systems (Fig. 7), here the oscillations are noisy and spatially homogeneous, without any apparent structures. Moreover, $M_{A}(t)$ and $M_{B}(t)$ display coherent oscillations with a fixed period and phase shift between the species.
Simulations with varying system sizes indicate there is a well-defined phase transition between a disordered phase with $R=\mathcal{L}=0$ to a swap phase with non-zero $R$ and $\mathcal{L}$. Figure 12 shows a color map of $R$ in the $(\tilde{K},\tilde{J})$ space, for $L=320$, while Fig. 13 shows a cut of the color map for $\tilde{K}=0.3$ with different $L$ values. Unlike in 2D (Fig. 10), here there is a critical $\tilde{J}$, below which $R$ approaches zero with $R\sim L^{-d/2}$ scaling and above which it approaches a finite non-zero value as $L$ increases, see inset of Fig. 13a. Moreover, the phase transition appears to be second-order in the sense that $R$ and $\mathcal{L}$ change continuously, and only their derivative is discontinuous.
[…]
The second-order phase transition is limited to small $\tilde{J}$ and $\tilde{K}$ values. This is seen in Fig. 12 where the change in $R$ and $\mathcal{L}$ at the phase transition appears to be smooth for small ($\tilde{J}$, $\tilde{K}$) values and discontinuous for higher values. In the next section, we focus on the second-order phase transition and discuss the other type in Sec. VIII.3.
Fig. 12: Order parameters in a large 3D system. Color map of (a) $R$ and (b) $\mathcal{L}$ as a function of $\tilde{J}$ and $\tilde{K}$ for a 3D system with $L=320$. In (a), red dot marks the pitchfork bifurcation, arrows corresponds to fixed $\tilde{K}$ and fixed $\tilde{J}$ cuts analyzed in Figs. 131416 and 20 and stars depict points analyzed in Fig. 15. Obtained from MC simulations.
Fig. 13: Phase transition from disorder to swap in 3D. (a) $R$ as a function of $\tilde{J}$ for $\tilde{K}=0.3$ shown for different linear system size $L$. Inset: $R$ vs. $L^{-1}$ for $\tilde{K}=0.3$ and $\tilde{J}=1.38,1.40,1.42,1.43,1.45$ from bottom to top on a log-log scale. (b) $\mathcal{L}$ as a function of $\tilde{J}$ with the same parameters as panel a. Obtained from MC simulations.
Fig. 14: Phase transition in 3D as a function of $\tilde{K}$. $R$ vs. $\tilde{K}$ for $\tilde{J}=1.5$ in a 3D system. Phase-space angular momentum is shown in the inset. At $\tilde{K}\approx 0.5$, a phase transition from swap to disorder occurs. Obtained from MC simulations.
VIII.1 Critical exponents
Having numerically confirmed the existence of a phase transition from disorder to swap in 3D, we move on to determine the associated critical exponents of $R$. Note that $\mathcal{L}$ is not expected to yield novel critical exponents because it is proportional to $R^{2}/T_{\rm osc}$ where $T_{\rm osc}$, the oscillation period, is finite at the Hopf bifurcation (see Sec. V.2.2).
As predicted by the MF analysis, the phase-space trajectories become circular when approaching the transition line (Fig. 15), supporting a CGL-like behavior as explained in Sec. V.2.2. Moreover, renormalization group studies based on the $\epsilon$-expansion suggest that the CGL equation with additive noise has the same critical exponents as the XY model [85, 86]. Thus, in what follows we will compare our results to the 3D XY critical exponents and to the 3D Ising critical exponents corresponding to the $\tilde{K}=0$ case.
Fig. 15: Phase-space trajectories close to and far from the transition from disorder to swap in 3D. $M_{B}$ as a function of $M_{A}$ in steady state for $\tilde{J}=1.35$ (cyan) and $\tilde{J}=1.5$ (magenta), with $\tilde{K}=0.1$ and $L=320$ (see points marked by stars in Fig. 12). Obtained from MC simulations.
According to the theory of finite-size scaling, in a finite system the order parameter $R$ and the susceptibility $\chi$, defined as $\chi=L^{d}\left(\langle s^{2}\rangle-\langle s\rangle^{2}\right)$, satisfy the following relations in the vicinity of the critical point $\tilde{J}_{c}$ (assuming fixed $\tilde{K}$[163],
$\begin{split}R&=L^{-\beta/\nu}\hat{R}(L^{1/\nu}(\tilde{J}-\tilde{J}_{c}))\\ \chi&=L^{\gamma/\nu}\hat{\chi}(L^{1/\nu}(\tilde{J}-\tilde{J}_{c})).\end{split}$ (52)
Here, $\hat{\chi}(x)$ and $\hat{R}(x)$ are called the scaling functions, and $\nu$, $\gamma$ and $\beta$ are the critical exponents $\xi\sim|\tilde{J}-\tilde{J}_{c}|^{-\nu}$ where $\xi$ is the correlation length, $\chi\sim|\tilde{J}-\tilde{J}_{c}|^{-\gamma}$, and $R\sim|\tilde{J}-\tilde{J}_{c}|^{\beta}$ (for $\tilde{J}>\tilde{J}_{c}$). The scaling functions are independent of $L$ and converge to non-zero constants at the critical point, i.e. $\hat{R}(0)={\rm const_{1}}$ and $\hat{\chi}(0)={\rm const_{2}}$. In what follows we assume that the finite-size scaling theory holds out of equilibrium, as is often done [164]. We set $\tilde{K}=0.1$ where numerical evidence shows that the transition is continuous (like a second-order transition).
The critical point $\tilde{J}_{c}$ is found using the Binder Cumulant [165] which is defined as $U_{L}=1-\langle s^{4}\rangle/(3\langle s^{2}\rangle^{2})$. Away from the critical point, the value of $U_{L}$ depends on system size, approaching $1/3$ in the disordered and $2/3$ in the ordered regime as $L\to\infty$ [166, 167]. However, at the critical point, in which correlation length exceeds the linear system size $L$, $U_{L}$ approaches a fixed value $U^{*}$ independent of $L$. Hence, the critical point is found from the intersection point of all the $U_{L}$ vs. $\tilde{J}$ curves plotted for different system sizes. Figure 16a shows that indeed all the curves cross at a single point ($\tilde{J}_{c}=1.34615\pm 0.00001$ for $\tilde{K}=0.1$).
Next, we plot the $\chi$ and $R$ as a function of $\tilde{J}$, see Fig. 16b-c. For each system size $L$, we find the maximum $\chi_{\rm max}(L)$ of $\chi$, the value $\tilde{J}_{\rm max}(L)$ of $\tilde{J}$ at the maximum, and the value $R(\tilde{J}=\tilde{J}_{c},L)$ of $R$ at the critical point. We then use the finite-size scaling relations (which can be derived from Eq. (52[163]),
$\begin{split}\tilde{J}_{\rm max}(L)-\tilde{J}_{\rm c}&=aL^{-1/\nu}\\ \chi_{\rm max}(L)&=bL^{\gamma/\nu}\\ R(\tilde{J}=\tilde{J}_{c},L)&=cL^{-\beta/\nu}\end{split}$ (53)
and perform linear fits by taking the logarithm of both equation sides (Fig. 16(d-f)). The procedure outlined here yields the critical exponents, where, to gain good accuracy, all averages are taken using $10^{6}$ independent samples of $s$. The error is estimated using the bootstrap method [163, 168]: the same procedure is repeated for multiple samples where each sample is constructed by drawing samples randomly, one at a time, from our data sample, and returning them to the data sample after they have been chosen. The standard deviation of the critical exponents is estimated by the standard deviation calculated from $10^{5}$ bootstrap resamplings of the data (Fig. 16(g-i)). Our final result is $\nu=0.68\pm 0.02$, $\gamma=1.33\pm 0.04$, and $\beta=0.35\pm 0.01$, which are in good agreement with 3D XY critical exponents $\nu_{XY}=0.67$, $\gamma_{XY}=1.32$, $\beta_{XY}=0.35$ [169, 170], more so than with 3D Ising critical exponents $\nu_{I}=0.63$, $\gamma_{I}=1.24$, $\beta_{I}=0.33$ [169], see graphical comparison in Fig. 17.
Fig. 16: Calculation of the critical exponents of the 3D NR Ising model. We set $\tilde{K}=0.1$. (a-c) The Binder cumulant $U_{L}$, the susceptibility $\chi$, and the order parameter $R$ as a function of $\tilde{J}$ for linear system sizes $L=16,20,24,28,32,36,40,44,48$ (from bottom to top in panel b). The vertical dashed grey line indicates the critical point, $\tilde{J}_{c}=1.34615$. Each point is calculated by averaging over $10^{6}$ independent samples of $s$. (d-f) $\chi_{\rm max}$, $|\tilde{J}_{\rm max}-\tilde{J}_{c}|$ and $R(\tilde{J}_{c})$ as a function of $L$ on a log-log scale. Dots are extracted from analyzing panels a-c and full lines indicate best linear fits. (g-i) Standard error of $\nu$, $\gamma$, and $\beta$ as a function of the number of bootstrap samples ($10^{5}$ bootstrap resamplings are performed in total).
Fig. 17: Compatibility with XY critical exponents. Critical exponents of the 3D NR Ising model (red) for $\tilde{K}=0.1$ extracted from finite-size scaling analysis along with their standard deviation represented by a semi-transparent red rectangle, and compared with the 3D Ising model (blue) and the 3D XY model (green).
VIII.2 Time-crystal behavior
On top of being globally ordered and time-dependent, the 3D swap phase possesses another notable property: its coherence time diverges because of its many-body nature, making it a time crystal [128, 134, 135, 136, 129, 137, 138].
To understand this feature, consider a single limit cycle oscillator subject to noise. This can be represented by the zero-dimensional MF equation of the NR Ising model with added noise, namely
$\partial_{t}m_{\alpha}=-m_{\alpha}+\tanh(\tilde{J}m_{\alpha}+\tilde{K}\varepsilon_{\alpha\beta}m_{\beta})+\eta_{\alpha}\;,$ (54)
where $\eta_{\alpha}$ is white noise satisfying $\langle\eta_{\alpha}(t)\rangle=0$ and $\langle\eta_{\alpha}\left(t\right)\eta_{\beta}\left(t^{\prime}\right)\rangle=\zeta\delta_{\alpha\beta}\delta\left(t-t^{\prime}\right)$ with $\zeta$ determining the noise amplitude. Figure 18 shows a numerical solution of Eq. (54) (panel a) along with the auto-correlation function of $m_{A}$ (panel d). Such auto-correlation of an observable $x$ is defined here as
$C_{x}\left(\tau\right)=\frac{\langle x\left(t\right)x\left(t+\tau\right)\rangle_{t,\Omega}}{\langle x\left(t\right)^{2}\rangle_{t,\Omega}}.$ (55)
Note that $\langle x\rangle=0$. If this was not the case, one should replace $x(t)$ with $x(t)-\langle x\rangle$ in Eq. (55). As seen in Figure 18, even though the magnetization performs quit regular oscillations, its auto-correlation function decays, thereby indicating that this oscillator eventually forgets its initial phase [171, 172, 124, 173, 174, 175, 176, 177].
In many-body systems, however, interactions can overcome noise, producing coherent oscillations over arbitrarily long periods [122, 95, 98, 108, 178, 110, 112, 94, 93, 111, 179, 106, 107, 110, 109, 180, 178, 96, 181, 182, 91, 90]. Figure 18 shows that in the 3D NR Ising model, temporal correlations persist and do not decay over large times (panels b-c and e-f), making it a robust clock. We stress that this robust clock is an emergent feature of our model which does not rely on any periodic drive externally imposing a phase.
The different behaviors of the time correlations in the noisy clock Eq. (54) and in the 3D NR Ising can be respectively compared to spatial density correlations in a liquid (which decay with distance) and a solid (which is maintained over arbitrarily large distances). Such comparison makes (i) a “time liquid” and (ii) a “time crystal” which has infinite coherence time.
We find the coherence time of the swap phase in the 3D NR Ising model by determining the decay rate of the auto-correlation’s envelope, $\tilde{C}_{M_{A}}(\tau)$. We construct this envelope $\tilde{C}_{M_{A}}(\tau)$ by interpolating the peaks of $|C_{M_{A}}(\tau)|$. We then implicitly define the coherence time $\tau_{c}$ as the time for which
$\tilde{C}_{M_{A}}\left(\tau_{c}\right)=\exp(-1).$ (56)
Figure 19 shows that $\tau_{c}$ increases with system size as
$\tau_{c}(L)\sim L^{3}\;,$ (57)
i.e. linearly with the number of spins, thereby proving the perfect time-crystal behavior of the 3D NR Ising in the thermodynamic limit. Note that a similar scaling for $\tau_{c}$ has been observed in a periodically driven 2D system in Ref. [183].
We can combine Eq. (57) with the conjecture of Ref. [125] to obtain a bound on the minimum rate of entropy production required to maintain the oscillations. This conjecture bounds the average entropy per oscillation $\Delta S$ by the number of coherent oscillations $\mathcal{N}\equiv\tau_{c}/T_{\rm osc}$ through the inequality $\Delta S\geq 4\pi^{2}\mathcal{N}$. As a consequence, the rate of entropy production per unit volume $\dot{s}\equiv\Delta S/(L^{d}T_{\rm osc})$ must satisfy
$\dot{s}\geq\frac{4\pi^{2}}{T_{\rm osc}^{2}}\,\frac{\tau_{c}(L)}{L^{d}}.$ (58)
Note that in the stable swap phase $T_{\rm osc}$ does not depend on $L$ (it converges to a constant as $L\to\infty$, see Fig. 26a in Sec. IX.1). As $\tau_{c}(L)\sim L^{d}$, this bound converges to a finite value when $L\to\infty$, meaning that there is a finite cost per oscillator to maintain the time-crystal state. Note that the fit in inset of Fig. 19 provides the prefactor for the example considered there.
Fig. 18: Time correlations in the NR Ising model. (a) Time evolution of $m_{A}$ in the zero-dimensional MF equation with added noise. (b) Time evolution of a single spin $\sigma^{A}_{i}$ in the 3D NR Ising model. (c) Time evolution of the total magnetization of $A$-spins, $M_{A}$, in the 3D NR Ising model. (d-f) Auto-correlation function of $m_{A}$, $\sigma^{A}_{i}$, and $M_{A}$, respectively (defined in Eq. (55)). Panels a and d are obtained from numerical integration of Eq. (54) with the parameters: $\tilde{J}=1.2$, $\tilde{K}=0.12$, $\zeta=0.0009$, $dt=0.01$. Panels b-c and e-f are obtained from MC simulation with the parameters: $\tilde{J}=1.6$, $\tilde{K}=0.1$, $L=80$.
Fig. 19: Diverging coherence time of the swap phase. The envelope $\tilde{C}_{M_{A}}(\tau)$ of the auto-correlation function of the total magnetization of $A$-spins in the 3D NR Ising model for linear system sizes $L=12,16,20,24,28,32$ (from bottom to top). Coupling parameters $\tilde{J}=1.6$ and $\tilde{K}=0.1$. The coherence time $\tau_{c}$ (Eq. (56)) is calculated from the intersection of $\tilde{C}_{M_{A}}(\tau)$ with $e^{-1}$ (dashed black line). Inset: $\tau_{c}$ as a function of $L$ on a log-log scale. Dots are calculated from the data in the main panel and line is best linear fit. The coherence time is proportional to $L^{d}$, supporting a time-crystal behavior. Obtained from MC simulations.
Finally, we note that ordered oscillations and infinite coherence time are not decoupled properties. It is plausible that a locally coupled system can maintain ordered oscillations over arbitrarily large distances only if all its parts adhere to the same prescribed phase, determined by the initial conditions.
VIII.3 The phase transition from disorder to swap in the high $(\tilde{J},\tilde{K})$ regime
Fig. 20: Destruction of the second-order phase transition due to scroll-waves in the high $\tilde{K}$ and high $\tilde{J}$ regime. (a) $R$ vs. $\tilde{K}$ for $\tilde{J}=2$. phase-space angular momentum is shown in the inset. (b) Snapshot of a 2D slice from a 3D simulation with the parameters $\tilde{J}=2$, $\tilde{K}=0.9$, $L=320$, marked by a black hollow circle on panel a. (c) The same snapshot coarse-grained over $10\times 10$ lattice sites. The snapshots show the destruction of order by scroll waves (appearing as spiral defects due to the 2D projection). Obtained from MC simulations.
At high $\tilde{J}$ and $\tilde{K}$ values, the transition from disorder to swap is different than for small $\tilde{J}$ and $\tilde{K}$ values. This is seen in Fig. 12 where the change in $R$ and $\mathcal{L}$ at the phase transition appears to become more abrupt in the upper right region, compared to the lower left region. We further analyze this in Fig. 20a by plotting the order parameters as a function of $\tilde{K}$, for fixed $\tilde{J}=2$, and various system sizes. The plots of $R$ and $\mathcal{L}$ vs. $\tilde{K}$ do not seem to converge to a second-order phase transition in which the order parameters are continuous. Instead, an abrupt change in $R$ and $\mathcal{L}$ emerges at high system sizes, suggesting a discontinuity as in first-order phase transitions. A clear critical point cannot be identified since convergence is not reached at $L=320$, the largest linear system size in our simulations (see $0.8\lesssim\tilde{K}\lesssim 1.1$ region in Fig. 20a).
The sharp decrease in synchronization is caused by the emergence of highly noisy scroll waves that destructs the order. This is seen in Fig. 20b showing a 2D snapshot taken from the 3D simulation. Scroll waves are the 3D analog of spiral waves, having line defects that are typically closed into rings instead of point defects [136, 184, 84, 185] (see also Movie 3 and Fig. 29 in Sec. X in which the scroll waves are more visible). They also have spiral cross sections, which is seen in the 2D-cut $\theta$ snapshot in Fig. 20b and in its coarse-grained version, Fig. 20c (the latter figure makes the defects more apparent by averaging out the noise). We discuss scroll waves in more detail in Sec. X.
The destruction of the swap phase by scroll waves is slightly reminiscent of the destruction of order in 2D due to spiral defects (compare Fig. 10 with Fig. 20a), except that here, it appears that the destruction only occurs for a limited range of high $\tilde{J}$ and $\tilde{K}$ values whereas smaller values are protected. Due to the size limitation of our simulations we cannot prove what portion of the 3D swap phase is destructed by the scroll waves or what kind of phase transition is associated with this destruction, although it appears to be first-order-like (in the sense that the order parameters are discontinuous).
VIII.4 Why is the swap stable in 3D?
Fig. 21: Continuous symmetry breaking of the swap phase. (a) Comparison of discrete and continuous symmetry breaking in equilibrium and out-of-equilibrium models. (b) $M_{B}$ vs. $M_{A}$ for $\tilde{K}=0.5$, $\tilde{J}=2$, and $L=80$. i,ii, and iii show increasing numbers of MC sweeps. Since the oscillation period is incommensurate, all phases are sampled in the infinite time limit. Hence, a continuous symmetry is broken. (c) Oscillation period of the magnetizations, $T_{\rm osc}$, in the 3D system as a function of $\tilde{J}$ for $\tilde{K}=0.3$ and $L=80$. $T_{\rm osc}$ is measured as twice the average time between two subsequent sign flips of $M_{A}$. The continuous dependence of $T_{\rm osc}$ on the coupling parameter indicates incommensurate oscillations. (d-e) comparison between domain walls in the Ising model (first row) and in the NR Ising model (second row). At $t=0$, the two systems were at steady state and then perturbed by a domain wall. The domain wall is created by flipping the magnetization (Ising) and shifting the oscillation phase (NR Ising) for all spins located at $y>L/2$. System parameters are: $\tilde{J}=1.4$ (Ising) and $\tilde{J}=1.5$, $\tilde{K}=0.1$ (NR Ising), and $L=320$. i and iii are Snapshots of $\sigma$ (Ising) and $\theta$ (NR Ising, see color code in Fig. 2) in a 2D slice taken from a 3D simulation at $t=0$ and $t=1157$, respectively. ii Kymograph of $|{\rm d}{\bar{m}}/{\rm d}{y}|$ (Ising) and $|{\rm d}{\bar{\theta}}/{\rm d}{y}|$ (NR Ising), defined by the discrete symmetric derivative ${\rm d}f/{\rm d}y\equiv\left|\left(f_{y+5}-f_{y-5}\right)/10\right|$, where $\bar{m}$ is the magnetization and $\bar{\theta}$ is the phase, both averaged over the $xz$ plane. The smearing out of the dark pink region indicates the dissipation of the domain wall. Panels b-e are obtained from MC simulations.
Our numerical evidence point to the existence of a stable swap phase. This relates to a longstanding question: can time-periodic states of noisy many-body systems with short-range interactions be stable in the thermodynamic limit [95, 98, 122, 108, 178, 110, 112, 94, 93, 111, 179, 106, 107, 110, 109, 180, 178, 96, 181, 182]. References [95, 98, 96] distinguished two cases, depending on whether the spontaneously broken time-translation invariance is discrete or continuous (Fig. 21a). In systems that evolve in discrete time as our MC simulation, this depends on the type of oscillations: commensurate vs. incommensurate. When the period of the oscillations is commensurate, i.e., a multiple of the discrete time-unite (a case referred to as “periodic $k$-cycle”), a discrete time-translation symmetry is broken. In that case, it was argued in [95] that the growth of out-of-phase droplets  [^4] Note that these potential droplets are not the droplets that destabilize the static-order phase (Fig. 23). Instead, these are droplets of a homogeneously oscillating phase, immersed in an oscillatory background destabilizes long-range ordered oscillations. That is because flat domain walls are expected to have a non-vanishing velocity (unlike equilibrium [187]), indicating that a fluctuation producing a large enough droplet, can expand rather than shrink (similar destabilization mechanism is observed in discrete-symmetry flocks [188]).
In the incommensurate case, i.e. when the oscillation period is not a multiple of the discrete time-unite (e.g. the time step of the MC simulation), the discrete-time system is equivalent to continuous-time one (the discreteness is irrelevant). The oscillatory state then breaks a continuous time-translation symmetry, which can be seen by the fact that the oscillation’s phase at a given time can take any value between 00 and $2\pi$. In continuous symmetry breaking, domain walls smear out (as in the XY model) and thus the destabilization-by-droplets argument does not apply [98, 96].
In the NR Ising model studied here, the oscillation period is incommensurate with the Monte-Carlo time step. This is seen both in Fig. 21b where all the points on the limit cycle are visited in the infinite-time limit, and in Fig. 21c where the period of the oscillation, measured as twice the average time between two subsequent sign flips of $M_{A}$, is shown to change continuously with the interaction parameters (a continuously changing variable is in general incommensurate with $1$). This means that, while stable oscillations are not guaranteed, they are not ruled out by the droplet argument in Ref. [95], see also Ref. [182]. Figure 21d-e further exemplifies this: it shows an analysis of a 3D simulation in which the system is perturbed from a stable swap state by introducing a flat domain wall. As time evolves, the domain wall is shown to widen and dissipate (on top of slightly moving in the $+\hat{y}$ direction), unlike a domain wall in the Ising model, supporting the picture of no steady domain-walls between oscillating states.
The broken continuous symmetry provides a reasoning for the resemblence of the NR Ising model with the XY ($O(2)$) model, observed both in the absence of time-dependent order in 2D and its stabilization in 3D with the corresponding critical exponents calculated in Sec. VIII.1. Moreover, it suggests an analogy with the Mermin-Wagner theorem [105, 81] in the time domain [104, 87]: continuous changes in the oscillation phase over space may behave as generalized Goldstone modes and their fluctuations preclude order in 2D yet allow for order in 3D and beyond.
An analogy between (de)synchronization and surface roughening provides further reasons to expect the swap phase to be stable in 3D and not in 2D. If we add random noise to the MF equation (28), it takes the form of a noisy reaction-diffusion system
$\frac{\partial\vec{X}}{\partial t}=F(\vec{X})+D\nabla^{2}\vec{X}+{\rm noise},$ (59)
where $\vec{X}=x_{1},..,x_{n}$ and $D$ can be a matrix. Now assume that the noiseless and diffusion-less version of Eq. (59) has a limit cycle in steady state. To test whether the addition of diffusion and noise allows for synchronized oscillations, we assume that we are deep within the synchronized regime, where fluctuations from the global limit cycle are small and can be treated perturbatively. In this case, we can focus on the phase variable that describes the position of the system on the limit cycle, while neglecting other degrees of freedom. This is known as a phase reduction, see Refs. [151, 189, 190, 191]. Then, Eq. (59) can be mapped to the known Kardar–Parisi–Zhang (KPZ) equation [151, 98, 101, 102],
$\frac{\partial\Psi}{\partial t}=\alpha\nabla^{2}\Psi+\beta\left(\nabla^{2}\Psi\right)+{\rm noise}$ (60)
where the deviation $\Psi$ of the local phase with respect to the phase of the deterministic limit cycle plays the role of the surface height in the KPZ equation (Fig. 22). The KPZ equation predicts that interfaces are always rough in 2D, but that there is a phase transition between rough and smooth interfaces in 3D [192]. This suggests, by analogy, that a coherent oscillating phase can exist in 3D, but not in 2D. Note, however, that the variable $\Psi$ in Eq. (60) is an angle (defined on a compact space, the circle $S^{1}$), so the results on the usual KPZ equation cannot be directly applied. In particular, the fact that the phase $\Psi$ is compact allows for the existence of topological defects, that may change the physics. We refer to Refs. [101, 102, 98, 193, 194, 195] for further details on the compact KPZ equation. In 3D, defect rings are expected to be small (at least for certain range of $\tilde{J}$ and $\tilde{K}$) and hence not to completely destroy long-range order [98].
Fig. 22: Mapping between surface roughening and desynchronisation. Schematic drawing of the mapping between coupled limit cycle oscillators (top) and surface roughening described by the KPZ equation (bottom). The phase $\varphi_{i}$ of the oscillator at position $i$ is mapped to a height $h_{i}$.
IX Destruction of static-order phase via droplets
In finite systems, a ferromagnetic-like state with static order is observed both in 2D and 3D. As system size increases, however, our numerical simulations indicate that this static state is destabilized. This is supported by Figs. 10 and 13 (right parts) showing that the $\tilde{J}$-transition-point between static order and oscillations in finite systems (indicated by a sharp change in the slope of $R$ and $\mathcal{L}$ becoming zero) increases with system size in a way that does not seem to converge to a finite $\tilde{J}$. Figures 14 and 20a showing $R$ and $\mathcal{L}$ as a function of $\tilde{K}$ have a very small region of static order, when $\tilde{K}$ is close to zero. However, this region too shrinks when $L$ increases.
The static-order phase is destabilized by nucleation of droplets that grow and flip the magnetization, in alternating order of A- and B-spins. This mechanism is demonstrated in Fig. 23, where snapshots of the spins at different times are shown for increasing system sizes, in a 2D system. The full evolution is shown in Movie 4. For $L=10$ (Fig. 23a), both species do not change their magnetization. There are attempts of B-spins to nucleate droplets, but those shrink and disappear. For $L=40$ (Fig. 23b), the system is no longer static: a handful of B-spins droplets (one clearly observed in the figure) surpass a certain critical droplet size, grow and flip the magnetization. Then, the same occurs for the A-spins, and so on, leading to an oscillatory behavior of $M_{A}(t)$ and $M_{B}(t)$, but with intermediate periods in which $M_{A}(t)$ and $M_{B}(t)$ are constants. When $L=200$ (Fig. 23c), multiple droplets ($\gg 1$) drive the magnetization flipping of each species, making the $M_{A}(t)$ and $M_{B}(t)$ curves smooth at all times, i.e, there is no finite period of time in which the magnetization is constant and the system could be called static. We note that the figure shows a 2D system but a similar behavior is observed in 3D.
Fig. 23: Instability of the static order phase in the thermodynamic limit due to droplet nucleation. Top four rows: Snapshots of $\sigma^{A}$ and $\sigma^{B}$ in a 2D system with linear system size (a) $L=10$, (b) $L=40$, (c) $L=200$. Time evolves from top to bottom. Fifth row: the corresponding total magnetizations, $M_{A}$ and $M_{B}$, as a function of time. Obtained from MC simulation with parameters $\tilde{J}=2.8$ and $\tilde{K}=0.3$, and with initialization at $t=0$ with all spins up.
To see why non-reciprocity allows droplets to grow, in contrast with the equilibrium Ising model where they tend to shrink, assume first that the system is in the static-order phase, so most spins in both lattices are up (Fig. 24a) [92]. Since the system is static, it can be mapped into two equilibrium Ising models with opposite magnetic fields, with energy
$E=-J\sum_{\langle i,j\rangle}\sigma_{i}\sigma_{j}-H\sum_{i}\sigma_{i}$ (61)
where $H\approx K$ for A-spins and $H\approx-K$ for B-spins (see Eq. (11) and Fig. 24b). While sub-system A is in a global free energy minimum in the effective equilibrium system, sub-system B is in a metastable state, as it prefers a state with opposite magnetization. Sub-system $B$ then transitions to its global minimum by nucleating droplets larger than a critical size $\rho_{c}$ above which the effective magnetic field, which acts on the volume of the droplet, overcomes the nearest neighbor attraction, imposing surface tension, which makes the droplet grow [196, 92] (Fig. 24c). After a droplet has expanded beyond $\rho_{c}$, the stable sub-system becomes metastable and nucleates droplets, and so on (Fig. 24d). Crucially, due to the symmetry between $A$ and $B$ spins, both droplet types expand with the same radial velocity. Therefore, a nested droplet remains much smaller than its “mother-droplet” until the latter finishes expanding (see Sec. XII in which this symmetry is absent). A simplified view of this mechanism can be found in defining a “selfish free energy” for the zero-dimensional MF equation, see Appendix B.
Fig. 24: Schematic drawing of the droplet argument for the instability of the static-order phase. (a) A system in the static-order phase can be reduced into (b) two equilibrium Ising models subject to magnetic fields, where one system is in a stable equilibrium while the other in a metastable state. (c) The metastable system nucleates droplets that expand and flip the magnetization. (d) The stable system becomes metastable and nucleates droplets that expand with the same velocity, and so on.
The argument sketched above is contingent upon $\rho_{c}$ being finite, which is the case in any finite dimension (and finite couplings $\tilde{J}$ and $\tilde{K}$). Hence, we conjecture that in the fully anti-symmetric NR Ising model, the static-order phase is unstable in any finite dimension.
IX.1 Between static-order and swap: the droplet regime
The droplet mechanism explained above means that as system size increases, a regime of droplet-induced swap develops instead of the static-order phase. In 2D, this regime separates static-order from disorder, while in 3D, it separates static-order from the noisy homogeneous swap regime.
In this regime, as $\tilde{J}$ increases (and as $\tilde{K}$ decreases), the critical droplet size increases, and as a results the droplets become more sparse and the distance they travel before colliding with each other increases. Eventually, $\tilde{J}$ becomes large enough (or $\tilde{K}$ small enough) that the critical droplet size becomes larger than the system size. In that case, the magnetization can flip when fluctuations are of the order of the system size, in the same way as in the usual equilibrium Ising model.
This is demonstrated in Fig. 25 where 2D-slice snapshots of the $\theta$ field, as well as the magnetizations of the two species as a function of time, are shown for a 3D system with linear size $L=80$ at increasing values of $\tilde{J}$. For small $\tilde{J}$, the 3D system is in a noisy homogeneous swap phase. There are no apparent droplets but rather noisy patterns without a well-defined shape (Fig. 25a,e). Put differently, independent fluctuations do not have time to expand significantly before colliding. For higher $\tilde{J}$ values we observe droplets that drive the magnetization flips (Fig. 25b-c,f-g). For an even larger $\tilde{J}$, no droplets are observed and the magnetizations do not change over a long period of time (Fig. 25d,h).
Fig. 25: Transition from swap to static order in a finite 3D system, as a function of $\tilde{J}$. Top: 2D-slice simulation snapshots of the $\theta$ field for varying values of $\tilde{J}$. Bottom: The corresponding total magnetizations $M_{A}$ (blue) and $M_{B}$ (red) as a function of time measured in simulation sweeps. (a,e) $\tilde{J}=1.6$. (b,f) $\tilde{J}=2.2$. (c,g) $\tilde{J}=2.3$. (d,h) $\tilde{J}=2.5$. Other parameters are $\tilde{K}=0.3$ and $L=80$. As $\tilde{J}$ increases, a noisy homogeneous swap phase is replaced by droplet-induced swap and then becomes static and ordered.
The droplet regime is characterized by a distinct behavior of the oscillation period as a function of system size. In Fig. 26a, the oscillation period in 3D, $T_{\rm osc}$, is shown as a function of $\tilde{J}$ and a fixed $\tilde{K}$, for increasing $L$ values. In the non-droplet regime, i.e. small $\tilde{J}$, $T_{\rm osc}$ is independent of system size while in the droplet-induced swap regime (large $\tilde{J}$) $T_{\rm osc}$ decreases with system size at intermediate $L$, until it converges in the $L\to\infty$ limit to a finite value, independent of $L$ (see overlap of orange and green lines in Fig. 26a up to $\tilde{J}\approx 2.3$). In the large $\tilde{J}$ limit, and for fixed system size, droplets are so rare that each magnetization flip is driven by a single droplet. In this regime, the droplet can emerge in any lattice position at a given time step, thus $T_{\rm osc}$ scales with $L^{-3}$ [196, 92] (see Fig. 26b).
The different oscillation regimes share similarities with the relaxation of the magnetization in an Ising model from an initial state magnetized opposite to the applied field $H$. Reference [196] identifies four regimes, depending on the field strength and system size. The “strong-field” regime in which “the droplet picture is inappropriate”, and three other regimes which are classified using droplet theory by the relation between three length scales: the critical droplet radius $\rho_{c}$, the mean droplet separation $\rho_{0}$, and the linear system size $L$. These are the multi-droplet regime $L\gg\rho_{0}\gg\rho_{c}$, the single-droplet regime $\rho_{0}\gg L\gg\rho_{c}$, and the “coexistence region” $\rho_{0}\gg\rho_{c}\gg L$ where the system behaves as if there is no applied field. The average metastable lifetime calculated in Ref. [196] is shown to be independent of system size in strong field and multi-droplet regime yet highly dependent on system size in the single-droplet and coexistence regime, in a way reminiscent of Fig. 26a. We note that the analogy drawn here is not perfect because non-reciprocal interactions are different from an externally applied field in that they change once the magnetization flips.
Fig. 26: Behavior of the oscillation period in the swap phase. (a) The oscillation period of the total magnetizations $T_{\rm osc}$ in the 3D system, on a logarithmic scale, as a function of $\tilde{J}$ for fixed $\tilde{K}=0.3$ and different system sizes. (b) The same with a rescaled oscillation period, $L^{3}T_{\rm osc}$. Dashed grey line in the top panel separates a regime in which the oscillations are driven by well-defined droplets (right) from a regime in which they do not (left). Obtained from MC simulations.
IX.2 The fate of the droplet regime
We now discuss the fate of the droplet-induced swap regime in the thermodynamic limit. In 2D, this regime is destabilized by spiral defects making it disordered, as was shown in Sec. VII.
In 3D, the scenario is more complex. While droplets persist at large system size, distant regions become less synchronized and the system loses its global order, see Fig. 27. While $R$ clearly decreases (see right region in Fig. 13), it does not converge with system size. Thus, it is unclear to what extent order is lost in the thermodynamic limit, and whether it remains finite or is ultimately deemed disordered ($R=0$). If the latter is correct, it remains an open question what is the type of phase transition that occurs and where lies the critical line between swap and disorder in this droplet regime. We remark that the decrease in order depicted in Fig. 27, suggests a coexistence of domains with different $\theta$ values (where $\theta$ is uniform within each domain), that may indicate a first-order-like transition. However, the domains are not static: they change over time due to droplet nucleation, growth, and collisions.
Fig. 27: Loss of synchronization in the droplet regime with increased system size. Snapshots of the $\theta$ field in 2D slices taken from a 3D MC simulation with $L=80,160$, and $320$. Other system parameters are $\tilde{K}=0.3$ and $\tilde{J}=2.25$. As system size increases, more $\theta$ values coexistence (seen as more colors in a single snapshot).
The decrease in order in the droplet-induced-swap regime causes an intriguing behavior, observed only for large system size, ($L\geq 160$): synchronization is non-monotonic as a function of $\tilde{J}$ (or $\tilde{K}$). See $L=320$ line in Fig. 13, where $R$ increases monotonically beyond the Hopf bifurcation until at $\tilde{J}\sim 2.05$, and then starts decreasing. It then reaches a local minimum at $\tilde{J}\sim 2.25$ and increases again until saturation at $\tilde{J}\sim 2.55$. A similar behavior is observed in the left part Fig. 20a, for $L=320$, where $R$ is non-monotonic as a function of $\tilde{K}$. In both cases the anomalous decrease starts when the swap phase transitions from noisy homogeneous swap to droplet-induced swap.
The mechanism under which global synchronization decreases with increased $\tilde{J}$ (or decreased $\tilde{K}$) in the droplet regime is shown in Fig. 28, which compares two simulations with the same $\tilde{K}=0.3$ value: one at the onset of droplet-induced swap ($\tilde{J}=2.04$), and the other well within the droplet-induced swap ($\tilde{J}=2.24$). Simulation snapshots (Figure 28a) reveal that while the system is more correlated over short distances at larger $\tilde{J}$ values, it loses global order due to large regions that are out-of-phase with the rest of the system (Fig. 28a). This can be quantified by a size dependent synchronization parameter,
$R(l)\equiv\left\langle s(l)\right\rangle\,\,\,\,\text{with}\,\,\,s(l)\equiv\frac{1}{l^{d}}\bigg{|}\sum_{j}{\rm e}^{i\theta_{j}}\bigg{|}$ (62)
where the sum is over all sites inside a box of linear size $l\leq L$, and the average is taken with respect to different boxes and time. The size dependent synchronization parameter is shown in Fig. 28b, for squares instead of cubes to match with the 2D snapshots. For small $l$, $R(l)$ is larger for the higher $\tilde{J}$ value while for large $l$, $R(l)$ is larger for the smaller $\tilde{J}$ value. Moreover, while $R(l)$ at low $\tilde{J}$ decays exponentially with $l$, in the high $\tilde{J}$ regime it does not and the second derivative of $R(l)$ changes signs. This means that within the droplet-induced swap regime, the system cannot be described by a single correlation length, and instead has multiple length scales related to the droplet parameters such as the critical droplet size, the typical distance between droplets, and the correlation length inside a single droplet.
Fig. 28: Decrease in global order in the droplet regime as a function of $\tilde{J}$. (a) Snapshots of 2D slices taken from a 3D simulation with $\tilde{K}=0.3$ and $L=400$, comparing two different $\tilde{J}$ values, $\tilde{J}=2.04$ (left) and $\tilde{J}=2.24$ (right). The top snapshots are taken shortly after initialization while the bottom snapshots are taken after 10000 MC sweeps. (b) The synchronization parameter $R(l)$ defined in Eq. (62), measured at steady state as a function of $l$, for the two $\tilde{J}$ values in panel a. Averages are taken over $(L/l)^{2}$ squares and 1000 MC sweeps.
X Coarsening dynamics: scroll waves and planar waves
We now qualitatively discuss the way in which the system coarsens into an ordered state in 3D. In the simulations used to calculate the order parameters $R$ and $\mathcal{L}$ (Figs. 12, 13,14, and 20), the spins were initialized in an ordered initial condition (all spins up), making the transition to steady state in the swap and static-order phases relatively fast. When fully disordered initial conditions are used (each spin is set to $+1$ or $-1$ randomly), however, the system can coarsen into long-lived metastable states, before reaching steady state. To show this, we present in Fig. 29a the same plot as Fig. 13a ($R$ vs. $\tilde{J}$ in 3D) but with random initial conditions, and average taken over a single simulation running for a very long time. In both the swap phase ($1.4\lesssim\tilde{J}\lesssim 2.3$) and static order phase ($\tilde{J}\gtrsim 2.3$) we see that for some of the realizations the system reaches the same steady state as with ordered initialization (shown in semi-transparent lines). However, sometimes the measured $R$ is much lower than the one obtained with ordered initialization, suggesting that the system got trapped in a metastable state. We cannot exclude that several equally stable states exist.
The states in which the system gets trapped in are scroll waves and planar waves, shown in Fig. 29b-d and Movie 3 in both 3D snapshots and 2D-slice snapshots. Here the scroll waves are much less noisy than those discussed in Sec. VIII.3.
[…]
The coarsening dynamics is most easily understood by following the defect filaments of the scroll waves. The defects are found using the size dependent synchronization order parameter, $s(l)$, in Eq. (62). Note that $s(l)$ is 1 if all the sites in the box are identical, and it is close to zero if the box has a mixture of sites in all four possible states. Thus, we classify defects as sites on the coarse-grained grid in which $s(l)$ is smaller than a threshold $s_{\rm max}$. The box size $l$ and the $s$-threshold are fine-tuned to detect the defect filaments while filtering out noise (in Fig. 29 $l=3$ and $s_{\rm max}=0.2$).
Figure. 29e shows an example of the dynamics of the defect filaments. At the beginning the system coarsens into a complex mixture of defect lines. After a while, only two defect lines survive. The lines are connected onto themselves by virtue of the periodic boundary conditions, having the topology of rings on a 3-dimensional torus with opposite winding numbers ($1$ and $-1$). Thus, they cannot shrink by themselves to a point. Finally, the two filaments collide, forming a new ring with zero winding number that shrinks and disappears. The system then stays in a planar waves state for a very long time. Movie 3 shows another example of the filament dynamics that ends in a defect-free swap phase.
The occurrences in which the system escapes scroll-waves state into swap or static order, without any observation of transitions from the swap or static order phase into scroll waves, suggest that these are metastable states. This is not the case, however, in the high $\tilde{K}$ limit, where the scroll waves appear to be stable, as explained in Sec. VIII.3.
[…]
XI Schematic phase diagram in the fully anti-symmetric case
In 2D, our numerical results suggest that both the swap phase and the static order phase are unstable, as long as $K\neq 0$. This leads to the conclusion that there is a single stable phase in the thermodynamic limit: disorder (see Fig. 30a). Note that the same conclusion applies to the 1D case, as is shown in Appendix D.
In 3D (see Fig. 30b), our numerical results suggest that static order is unstable but that the swap phase is stable, with a second-order-like phase transition from disorder to swap featuring 3D XY critical exponents. Regions with high $\tilde{J}$ or $\tilde{K}$ have not been revealed in full: in the large $\tilde{J}$ limit, synchronization decreases, possibly to zero, due to the difficulty of droplet-induced swap to synchronize. In the high $\tilde{K}$ regime, scroll waves completely destabilize the long-range order. Both mechanisms are capable of destroying a large portion of the stable swap phase regime, and there is also a possibility that they would meet in the high $\tilde{J}$ and high $\tilde{K}$ regime, leading to a compact swap phase in $(\tilde{J},\tilde{K})$ space rather than an infinite one. Due to the limitations of our simulations, these interesting possibilities lie beyond the scope of this work, which has focused mostly on the behavior at and close to the Hopf bifurcation, where the swap phase appears to be stable and critical exponents can be calculated.
Fig. 30: Schematic phase diagram of the fully anti-symmetric case. (a) Schematic 2D phase diagram and (b) schematic 3D phase diagram. To be compared with the MF prediction, Fig. 4. Green line is second-order-like phase transition with critical exponents compatible with those of the 3D XY model. Grey dashed lines are phase transitions that have not been revealed in full and are possibly first-order-like.
XII General asymmetric couplings
In the previous sections, we made two assumptions. First, we assumed that the interaction between the two spin species is fully anti-symmetric, i.e. $K_{+}=0$ in Eq. 12. Second, we assumed that both species have the same nearest-neighbors interaction strength $J$. These two conjoint assumptions lead to a symmetry between the two species which renders the model $C_{4}$ symmetric (instead of the less restrictive $\mathbb{Z}_{2}$ symmetry with only up-down symmetry). Therefore, we need to test whether our results rely on this symmetry or are robust to its breakdown.
In this section, we break the $C_{4}$ symmetry by considering a scenario where $K_{+}\neq 0$.
XII.1 Mean-field analysis
Expanding Eq. (23) in the small $m_{\alpha}$ and $\nabla m_{\alpha}$ limit, we obtain,
$\partial_{t}m_{\alpha}=$ $-(1-\tilde{J})m_{\alpha}+\tilde{K}_{\alpha\beta}m_{\beta}+D\nabla^{2}m_{\alpha}$ (63)
$-\frac{1}{3}\left(\tilde{J}m_{\alpha}+\tilde{K}_{\alpha\beta}m_{\beta}\right)^{3}.$
Taking $D=0$ (as was done in Sec. V.2 for the fully anti-symmetric case), the eigenvalues $\lambda_{1}$ and $\lambda_{2}$ of the Jacobian matrix $\mathcal{J}$ at $m_{A}=m_{B}=0$ are
$\lambda_{1,2}=-1+\tilde{J}\pm\sqrt{\tilde{K}_{+}^{2}-\tilde{K}_{-}^{2}}.$ (64)
Thus, upon changing $\tilde{J}$, the possible phases and bifurcations depend on the value of $\tilde{K}_{+}^{2}-\tilde{K}_{-}^{2}$, as shown in Fig. 31 for the the full homogeneous MF equation, Eq. (23) with $D=0$.
For $\tilde{K}_{+}^{2}-\tilde{K}_{-}^{2}<0$ ($\tilde{K}_{+}<\tilde{K}_{-}$) the system is disordered when $\tilde{J}<1$, and has a Hopf bifurcation at $\tilde{J}=1$, taking the system from disorder to a limit cycle (Fig. 31a). Note that unlike the $K_{+}=0$ case studied throughout most of the paper, here the limit cycle is not rotationally-symmetric even in the $\tilde{J}-1\ll 0$ limit. This is easily seen when writing the homogeneous MF equation in polar coordinate
$\begin{split}\partial_{t}r&=-\left(1-\tilde{J}-\tilde{K}_{+}\sin(2\theta)\right)r+\mathcal{O}\left(r^{3}\right)\\ \partial_{t}\theta&=-\tilde{K}_{-}+\tilde{K}_{+}\cos(2\theta)+\mathcal{O}\left(r^{2}\right)\end{split}$ (65)
where the leading order term in $r$ is $\theta$-dependent and a separation between the time scales of $r$ and $\theta$ can not be invoked as in Sec. V.2. Instead, the limit cycle is elliptical close to the Hopf bifurcation [151].
Fig. 31: Bifurcations in the general asymmetric case. Flow in $\left(m_{A},m_{B}\right)$ space calculated from the full MF equation, Eq. (23) with $D=0$ and $\tilde{K}_{-}=0.2$, and with (a) $\tilde{K}_{+}=0.1$, (b) $\tilde{K}_{+}=0.3$, and (c) $\tilde{K}_{+}=1.1$. In each panel, $\tilde{J}$ is increased from left to right to capture the different phases. In (a) $\tilde{J}=0.4,1.2,1.5,1.7$, in (b) $\tilde{J}=0.4,0.9,1.4,2$ and in (c) $\tilde{J}=0.1,2.3,3.1$. The arrows mark the bifurcations that connect every two configurations: Hopf, saddle-node on an invariant circle (SNIC), saddle-node (SN) or pitchfork (PF). The bifurcations are consistent with the analysis of the expanded MF equation (Sec. XII.1).
Increasing $\tilde{J}$ leads to additional bifurcations (Fig. 31a): first, a SNIC bifurcation where the limit cycle is destroyed by the emergence of two (rather than 4 for the $K_{+}=0$ case) pairs of stable and saddle points, and for an even larger $\tilde{J}$, a saddle-node bifurcation where two additional pairs of stable and saddle points emerge. We note that this behavior is observed for $\tilde{K}_{+}$ and $\tilde{K}_{-}$ much smaller than $1$ ($0.1$ and $0.2$ respectively in the Fig. 31). For larger values, a saddle-node bifurcation can occur such that the ferromagnetic phase coexists with the limit cycle (this scenario is explored in  [143]).
For $10\leq\tilde{K}_{+}^{2}-\tilde{K}_{-}^{2}<$ 1 the system is disordered when $\tilde{J}<1-\sqrt{\tilde{K}_{+}^{2}-\tilde{K}_{-}^{2}}$ and has a pitchfork bifurcation taking the system from disorder to static order with two stable fixed points. In that case, there is no oscillatory phase (Fig. 31b). Finally, for $\tilde{K}_{+}^{2}-\tilde{K}_{-}^{2}>1$ there are two stable fixed points already at $\tilde{J}=0$ (Fig. 31c) and thus the only stable homogeneous phase is the static order phase (although $\tilde{J}$ must be larger than zero for coarse-graining to take place). We note that the absence of oscillations for $\tilde{K}_{+}>\tilde{K}_{-}$ is due to the fact that the two species, while having non-reciprocal interactions, “agree” with each other on whether they like to align or anti-align, making a magnetization flip unfavorable for either species.
Figure 32 shows a cut through the three-dimensional (depending on $\tilde{J},\tilde{K}_{+}$, and $,\tilde{K}_{-}$) phase diagram, for a fixed $\tilde{K}_{+}$ (set to be $0.3$). The main difference between this phase diagram and the one in Fig. 4 showing the $\tilde{K}_{+}=0$ case, is the pitchfork bifurcation line from a paramagnetic to ferromagnetic phase in the $\tilde{K}_{-}\leq\tilde{K}_{+}$ region, which is merely a point in Fig. 4.
Fig. 32: Mean-field phase diagram of the general asymmetric case. The diagram is shown as a function of $\tilde{J}$ and $\tilde{K}_{-}$. Reciprocal inter-species interaction coupling is set to be $\tilde{K}_{+}=0.3$. The lines color code is the same as Fig. 4. Obtained from numerical analysis of Eq. (23) with $D=0$.
XII.2 Swap phase
We perform MC simulations in both 2D and 3D for the general asymmetric case, i.e. with the inclusion of reciprocal inter-species interactions. In 2D, we find that the swap phase is unstable, as in the case of fully anti-symmetric interactions (Sec. VII). This is shown in Fig. 33 in which $R$ and $\mathcal{L}$ are plotted as a function of $\tilde{J}$ with fixed $\tilde{K}_{+}$ and $\tilde{K}_{-}$. In the region in which $\mathcal{L}\neq 0$ (indicating the swap phase), both $R$ and $\mathcal{L}$ decrease with system size. In appendix E we show color maps of $R$ and $\mathcal{L}$ as a function of $\tilde{J}$ and $\tilde{K}_{-}$ with fixed non-zero $\tilde{K}_{+}$. The color maps show that, similarly to the $\tilde{K}_{+}=0$ case in Fig. 9, $R$ and $\mathcal{L}$ diminish as system size increases.
[…]
In 3D, MC simulations suggest that a stable swap phase still exists, even when reciprocal interactions are added. Figure 34a-b shows color maps of $R$ and $\mathcal{L}$ for a fixed $\tilde{K}_{+}$ and $L=80$. Figure 34c-d then analyzes a cut through the color map for fixed $\tilde{K}_{-}$. As $\tilde{J}$ is varied, a phase transition from disorder to swap is observed and becomes more sharp as $L$ increases. The swap phase is stable for $\tilde{J}\lesssim 1.7$ but destabilizes at higher values, as is discussed next.
To conclude, the instability of the swap phase in 2D and its stability in 3D are not limited to the case of fully anti-symmetric interactions in which $\tilde{K}_{+}=0$, but extends to the general asymmetric case.
Fig. 34: Stable swap phase in 3D in the general asymmetric case. (a) $R$ and (b) $\mathcal{L}$ color map as a function of $\tilde{J}$ and $\tilde{K}_{-}$ with $\tilde{K}_{+}=0.1$ and $L=80$. Dashed grey line shows the fixed $\tilde{K}_{-}$ cut analyzed in panels b and c. (c) $R$ and (d) $\mathcal{L}$ as a function $\tilde{J}$ with $\tilde{K}_{+}=0.1$ and $\tilde{K}_{-}=0.3$, for different system sizes. Inset in c shows a snapshot of $\theta$ in a 2D-cut of the 3D system, with $\tilde{J}=2$, $L=80$, and the same $\tilde{K}_{+}$ and $\tilde{K}_{-}$ values as in the main panel. Obtained from MC simulations.
XII.3 Stability of the static-order phase: a droplet-capture mechanism
While the fate of the swap phase is robust to the addition of reciprocal inter-species interactions, our MC simulations suggest this is not the case for the static-order phase.
First, let us look at the right hand side of Figs. 33 and 34c-d, showing the transition to static order ($R\neq 0$ and $\mathcal{L}\approx 0$) for a finite-size system in 2D and 3D, respectively. In 2D the $\tilde{J}$-transition-point between the unstable swap regime and static order increases with system size and it is hard to conclude whether static order is stable in the thermodynamic limit. In 3D, there is some indication that there is a well-defined phase transition from disorder to static order: for linear system sizes $L=40,80,160$, there is an abrupt transition from a region in which $R$ and $\mathcal{L}$ approach zero (see $1.8\lesssim\tilde{J}\lesssim 2.2$ in Fig. 34c-d) to a regime in which $R$ is finite and $\mathcal{L}\to 0$ ($\tilde{J}\gtrsim 2.2$ in Fig. 34c-d), and the transition seems to converge at $\tilde{J}\approx 2.2$. Surprisingly, the supposed phase transition occurs even though $\tilde{K}_{+}$ is lower than $\tilde{K}_{-}$ ($\tilde{K}_{+}=0.1$ while $\tilde{K}_{-}=0.3$), which is the case where one species favors alignment while the other favors anti-alignment. This begs the question: does the droplet destabilization argument sketched in Sec. IX for the fully anti-symmetric case apply here?
Let us assume again a static-order scenario in which both spin species are up on average. The species favoring anti-alignment ($B$) prefers a state in which all of its spins are flipped. As a result, it will nucleate droplets with opposite magnetization. As in the fully anti-symmetric case, we expect that once these droplets will surpass a critical size, they will expand instead of shrinking. After certain amount of time, the same event will repeat itself for the $A$ species: a critical droplet will nucleate and expand.
However, because of the asymmetry between the two species imposed by $K_{+}$, the critical droplet size $\rho_{c}$ and more importantly the velocity in which droplets expand $v(\rho)$, is different for $A$-droplets and $B$-droplets. Here, $A$-spins interact more strongly with $B$-spins than the other way around, leading to smaller $\rho_{c}$ and larger $v(\rho)$ for $A$-droplets [196]. As a result, a nested $A$-droplet inside a $B$-droplet will expand faster, eventually catching up with the boundaries of the $B$-droplets. Once this happens, our simulations show that the $B$-droplet can disintegrate, and the magnetization flip it prevented. Movie 5 shows a 2D simulation in which this mechanism takes place. Snapshots from the simulation are shown in Fig. 35.
Fig. 35: Droplet capture mechanism. MC simulation snapshots of a 2D system with $\tilde{J}=3.3$, $\tilde{K}_{+}=0.1$, $\tilde{K}_{-}=0.3$, and $L=150$. Each row displays a distinct time step, with time ascending from top to bottom. Species $B$ nucleates a droplet, but before it reaches system size, a nested droplet of $A$-spins nucleates and expands at faster speed, eventually catching up with the boundaries of the $B$-droplet, making it shrink and disappear. This mechanism can stabilize the static-order phase when $\tilde{K}_{+}<\tilde{K}_{-}$.
To summarize, when inter-species interactions are not fully anti-symmetric, there is a competing mechanism that prevents droplets from growing. This suggests a re-stabilization of the static-order phase in the general asymmetric case, for $d\geq 2$ (even if the numerical evidence from the order parameter plots is not conclusive).
XII.4 Absence of infinite-period phase transition
[…]
XII.5 Schematic phase diagrams in the general asymmetric case
A schematic drawing of the 2D and 3D phase diagram with non-zero $\tilde{K}_{+}$ is shown in Fig. 36. In 2D, we only find disorder and static-order phases, while in 3D, we observe all the phases predicted by mean-field: disorder, swap, and static order. We expect the disorder-to-swap phase transition to be characterized by the same critical exponents as in Sec. VIII.1 (compatible with 3D XY), at least throughout a portion of the phase transition line. Other portions may be described by first-order-like phase transitions. Similarly, we hypothesize that the phase transition from disorder to static order in both 2D and 3D has Ising critical exponents in the regime where $K_{+}>K_{-}$, but might be first-order-like in the opposite $K_{+}<K_{-}$ regime (see Fig. 34 in which the $K_{+}<K_{-}$ phase transition seems more abrupt than the $K_{+}>K_{-}$ phase transition). Finally, we cannot exclude that the three phases in 3D meet at a tricritical point or line in the three-dimensional phase diagram.
Fig. 36: Schematic phase diagram of the general asymmetric case in (a) 2D and (b) 3D, shown as a function of $\tilde{J}$ and $\tilde{K}_{-}$, with fixed (non-zero) $\tilde{K}_{+}$. The nature of the phase transition lines are left unspecified, see discussion in Sec. XII.5. The question mark represents the uncertainty about whether the swap and static order phases meet at a point or are strictly separated by disorder.
XIII Conclusions
We introduced a non-reciprocal Ising model as a minimal model for non-reciprocal many-body dynamics, and studied the stability of its static and time-dependent phases as well as the transitions between them. Our main result is the stability of a time-dependent (swap) phase in 3D, which behaves as a time-crystal with infinite coherence time in the thermodynamic limit. The phase transition from disorder to swap is characterized by critical exponents compatible with the 3D XY model, which can be rationalized by the breaking of a continuous symmetry (time translation) in the swap phase. The existence of a swap phase in 3D is robust, valid for fully anti-symmetric interactions between the species as well as in the more general case of asymmetric interactions.
Via simulations and theoretical arguments, we conjecture the static-order (ferromagnetic) phase to be unstable in any finite dimension when interactions are fully anti-symmetric. The instability is caused by the expansion of droplets of opposite magnetization occurring in the metastable species. When the inter-species interaction is not fully anti-symmetric, a competing mechanism of droplet-capture by the other species can re-stabilize the static-order phase. A direct phase transition between a time-dependent phase and a static-ordered one appears to be unfeasible.
Acknowledgements.
We thank G. Biroli, S. Diehl, O. Granek, M. Han, T. Khain, P. Littlewood, R. Mandal, D. Mukamel, S. Sethi, S. Sondhi, G. A. Weiderpass, C. Weiss, and T. Witten for helpful discussions. Y.A., D.S. and M.F. acknowledge support from a MRSEC-funded Kadanoff–Rice fellowship and the University of Chicago Materials Research Science and Engineering Center, which is funded by the National Science Foundation under award no. DMR-2011854. Y.A. acknowledges support from the Zuckerman STEM Leadership Program. D.M., M.F., and V.V acknowledge support from the France Chicago center through a FACCTS grant. M.F. acknowledges support from the National Science Foundation under grant no. DMR-2118415 and the Simons Foundation. V.V. acknowledges partial support from the Army Research Office under grant nos. W911NF-22-2-0109 and W911NF-23-1-0212 and the Theory in Biology program of the Chan Zuckerberg Initiative. This research was partly supported by the National Science Foundation through the Center for Living Systems (grant no. 2317138) and the National Institute for Theory and Mathematics in Biology (NITMB). All the authors acknowledge the support of the UChicago Research Computing Center which provided the computing resources for this work.
Appendix A The effect of different update rules
Out of equilibrium, different update rules may change the steady state of the system. To study this effect in the NR Ising model, we performed simulations with random sequential updates, checkerboard updates [152], and synchronous updates (in which all the spins are updated at the same time) [199]. Additionally, we used the Metropolis algorithm [163] and compared it to the Glauber which was used throughout the paper.
For fixed coupling parameters ($\tilde{J}$ and $\tilde{K}$), the different update rules changed the dynamics at steady state. However, they did not lead to qualitative changes in the phase diagram. That way, the critical point $\tilde{J}_{c}$, for some fixed $\tilde{K}$, is altered, but not the existence / destabilization of a phase transition. As an example, we show in Fig. 37 that the phase transition from disorder to swap in 3D, with random sequential updates (standard Glauber algorithm), is qualitatively the same as with the checkerboard algorithnm which was used to produced Fig. 13.
We did not find any qualitative difference related to phase transitions. The only notable difference is that with synchronous updates, fluctuations were considerably larger due to “checkerboard patterns” (small regions in which each spin is anti-aligned with its same-species neighbours) that frequently appeared and then disappeared.
Fig. 37: Qualitative agreement between two Monte-Carlo update rules. $R$ as a function of $\tilde{J}$ for $\tilde{K}=0.3$ and different linear system size $L$, calculated from MC simulations with random spin update rule instead of the checkerboard update which is used in Fig. 13a. Red semi-transparent line shows the checkerboard update results for $L=160$. Results are qualitatively the same as in Fig. 13a although exact $\tilde{J}_{c}$ value varies. Averages are taken over a smaller number of realizations than in Fig. 13a, making the high $\tilde{J}$ data noisy.
Appendix B Selfish free energy
The dynamics of non-conserved fields $m_{\alpha}(\vec{r},t)$ with $\alpha=1,...,N$ (model A) in equilibrium is [200, 27, 201]
$\partial_{t}m_{\alpha}(\vec{r},t)=-\mu_{\alpha}\frac{\delta F[m_{1},...,m_{N}]}{\delta m_{\alpha}(\vec{r},t)}+\eta_{\alpha}(\vec{r},t)$ (66)
where $F$ is the effective free energy functional, $\mu_{\alpha}$ are generalized mobilities, and $\eta_{\alpha}$ is a white noise satisfying
[…]
where $D_{\alpha}=\mu_{\alpha}k_{B}T$.
For non-reciprocal systems, there is no global $F$ generating the dynamics. However, we extend the notion of a free energy by writing the dynamics of the two magnetization fields $m_{A}$ and $m_{B}$ as
$\partial_{t}m_{\alpha}(\vec{r},t)=-\mu_{\alpha}\frac{\delta F_{\alpha}[m_{A},m_{B}]}{\delta m_{\alpha}(\vec{r},t)}+\eta_{\alpha}(\vec{r},t)$ (68)
where $F_{\alpha}$ is the selfish free energy of the $\alpha$-species. We set $\mu_{\alpha}=1$ and use Eq. (28) to obtain the selfish free energies of the NR Ising model in the fully anti-symmetric case,
$F_{A}$ $=\frac{1}{2}\left(1-\tilde{J}\right)m_{A}^{2}-\tilde{K}m_{A}m_{B}+D\left(\nabla m_{A}\right)^{2}$
$+\frac{1}{12\tilde{J}}\left(\tilde{J}m_{A}+\tilde{K}m_{B}\right)^{4}$ (69)
$F_{B}$ $=\frac{1}{2}\left(1-\tilde{J}\right)m_{B}^{2}+\tilde{K}m_{A}m_{B}+D\left(\nabla m_{B}\right)^{2}$
$+\frac{1}{12\tilde{J}}\left(\tilde{J}m_{B}-\tilde{K}m_{A}\right)^{4}.$
Equation (B) can be viewed as the free-energy landscape that each species “sees”, where the free-energy landscape of $A$ changes with $m_{B}$ and vice versa.
We now restrict ourselves to the zero-dimensional case ($D=0$), for simplicity. In order to see the free-energy landscape of each species, we solve the MF equation (Eq. (28) with $D=0$) and substitute the $m_{B}(t)$ solution into $F_{A}$ and the $m_{A}(t)$ solution into $F_{B}$. The free-energy landscape in steady state is shown in Fig. 38, for the disordered, swap, and static-order phases. The actual values of $m_{A}$ and $m_{B}$ in steady state are shown as disks moving on the free-energy landscape.
Fig. 38: Selfish free-energy landscape. $F_{A}$ as a function of $m_{A}$ (blue) and $F_{B}$ as a function of $m_{B}$ (red) in zero-dimension NR Ising model. $m_{A}(t)$ and $m_{B}(t)$, obtained numerically by solving Eq. (28) with $D=0$ at steady state, are substituted into $F_{A}$ and $F_{B}$ defined in Eq. (B), respectively. Actual $m_{A}$ and $m_{B}$ values are shown as disks. (a) Disorder, $\tilde{J}=0.9$, (b) swap, $\tilde{J}=1.11$, and (c) static-order, $\tilde{J}=1.3$, and in all plots $\tilde{K}=0.1$. In panel a, $A$ and $B$ lines as well as disks are identical and therefore only B is shown.
In the disordered phase (Fig. 38a), both $m_{A}$ and $m_{B}$ are in the global minimum of their respective free energies. In the swap phase (Fig. 38b), the free-energy landscape of each species changes continuously. While one species is momentarily in a global minimum, the other species is descending along the free-energy landscape which changes the free-energy landscape of the first species and so on. In the static order phase (Fig. 38c), one of the species is at its global minimum while the second is in a local minimum, that is not the global one. This configuration is stable in MF, but once noise is introduced, it becomes metastable: the second species will “jump” over the barrier and reach global minimum, making the first species metastable and the cycle repeats itself. In a spatially extended system, this transition happens via droplet nucleation and growth as depicted in Sec. IX.
Appendix C Simulations of the MF equation in 2D
Fig. 39: The effect of added noise on the MF equation in 2D. Simulations of Eq. (70) in 2D comparing the cases of no added noise with added noise and ordered vs. disordered initial conditions. First column is kymographs of the $\theta$ field for a single row on the 2D system. Second, third, and forth columns are snapshots of the $m_{A}$, $m_{B}$, and $\theta$ fields, respectively, taken after the system has reached steady state. Without added noise, the steady state is a homogeneous swap phase for ordered initial conditions and an inhomogeneous spiral phase for disordered initial conditions. With enough added noise, the swap phase is unstable even with ordered initial conditions and spiral defects are formed for both initialization types. Simulation time step is $\Delta t=0.05$, noise amplitude for the case of added noise is $\zeta=0.36$, and other system parameters are: $\tilde{J}=2$, $\tilde{K}=0.8$, $L=200$, and $\tau=1$. At $t=0$, we set $m_{A},m_{B}=1\pm 0.1$ for ordered initial conditions and $m_{A},m_{B}=\pm 0.1$ for disordered initial conditions, where the plus/minus sign is chosen at random for each spin. Kymographs show evolution between $t=9900$ and $t=10000$.
We simulate the full MF equation on a grid (Eq. (21)) in 2D in the regime corresponding to a stable swap phase, and in the fully anti-symmetric case. Results are shown in Fig. 39. Depending on initial conditions, the system can end up in different states. An initially ordered state leads to a stable swap phase where the $\theta$ field changes periodically and homogeneously (Fig. 39, first row), while an initially fully disordered state ($m_{A}$ and $m_{B}$ are chosen randomly at each point to be either $+0.1$ or $-0.1$) reaches a stable state with spiral defects, whose cores barely move over time (Fig. 39, second row).
We test the effect of adding noise to the MF equation by writing,
$\partial_{t}m_{i}^{\alpha}=-\frac{m_{i}^{\alpha}}{\tau}+\frac{1}{\tau}\tanh\left(\frac{J\sum_{j\,{\rm nn\,of}\,i}m_{j}^{\alpha}+K\varepsilon_{\alpha\beta}m_{i}^{\beta}}{k_{B}T}\right)+\eta_{i}^{\alpha}.$ (70)
where $\eta_{i}^{\alpha}$ is white noise satisfying $\langle\eta_{i}^{\alpha}\rangle=0$ and $\langle\eta_{i}^{\alpha}\left(t\right)\eta_{j}^{\beta}\left(t^{\prime}\right)\rangle=\zeta\delta_{\alpha\beta}\delta_{ij}\delta\left(t-t^{\prime}\right)$. Above some critical $\zeta$, and for fixed system size, we observed that even if the system is initialized in an ordered state, after a long time, spiral defects form (see Fig. 39, third row): they are qualitatively similar to those observed when random initial conditions are implemented (Fig. 39, fourth row). In that sense, the MF equation with noise exhibits the same qualitative behavior as the Monte-Carlo simulations, where the swap phase is unstable in 2D.
Appendix D 1D NR Ising model
Fig. 40: 1D NR Ising model. (a) $R$ and (b) $\mathcal{L}$ as a function of $\tilde{J}$ for $\tilde{K}=0.3$ and different system size $L$. (c) Kymographs of $\theta$ for $\tilde{J}=3$ (left) and $\tilde{J}=4$ (right), comparing the Ising model ($\tilde{K}=0$, top) with the NR Ising model ($\tilde{K}=0.2$, middle and $\tilde{K}=0.5$, bottom). System size is $L=500$. Obtained from MC simulations.
We simulate the one-dimensional NR Ising model, with fully anti-symmetric interactions, using random picking Glauber dynamics as in the 2D simulations (see Sec. VI). The order parameters $R$ and $\mathcal{L}$ as a function of $\tilde{J}$ and for fixed $\tilde{K}$ are shown in Fig. 40a and Fig. 40b, respectively, while kymographs of $\theta$ with different $\tilde{J}$ and $\tilde{K}$ values are shown in Fig. 40c.
The NR Ising model does not have a phase transition to an ordered state (either static or time dependent), as is evident from the decrease of $R$ and $\mathcal{L}$ as system size increases. Similarly to the 1D Ising model, as $L$ increases, domains of opposite magnetization emerge and the system becomes disordered. However, in contrary to the Ising model, the boundaries between the domains do not diffuse but move with non-zero velocity (see Fig. 40c where the kymograph slope indicates the velocity). Domains expand in a way that can be viewed as the 1D version of droplet growth discussed in Sec. IX.
Appendix E 2D numerical phase diagram in the general asymmetric case
Figure 41 shows color maps of $R$ and $\mathcal{L}$ as a function of $\tilde{J}$ and $\tilde{K}_{-}$ with fixed non-zero $\tilde{K}_{+}$, for a 2D system with $L=20$ (panels a-b) and $L=80$ (panels c-d). While for $L=20$ the phase diagram is similar to the MF prediction (Fig. 32), at $L=80$ both $R$ and $\mathcal{L}$ diminish and there is no apparent swap phase, supporting the conclusions of Sec. XII.2.
Fig. 41: 2D numerical phase diagram in the general asymmetric case with fixed (non-zero) $\tilde{K}_{+}$. (a) $R$ and (b) $\mathcal{L}$ color maps as a function of $\tilde{J}$ and $\tilde{K}_{-}$, with $\tilde{K}_{+}=0.1$ and $L=20$. (c-d) The same with $L=80$. Obtained from MC simulations.
Appendix F Additional cuts through the 3D numerical phase diagram in the general asymmetric case
In the general asymmetric case, the phase diagram is three-dimensional as it depends on $\tilde{J}$, $\tilde{K}_{+}$, and $\tilde{K}_{-}$. In the main text the order parameters were calculated as a function of $\tilde{J}$ and $\tilde{K}_{-}$ for a fixed value of $\tilde{K}_{+}$ (Fig. 34). Figure 42 shows additional cuts, both for fixed $\tilde{J}$ and for fixed $\tilde{K}_{-}$, for a system of linear size $L=80$. In both cuts, the three phases of disorder ($R=0$ and $\mathcal{L}=0$), swap ($R\neq 0$ and $\mathcal{L}\neq 0$) and static-order ($R\neq 0$ and $\mathcal{L}=0$) are clearly shown. However, the transition between the swap phase and the static-order phase is mediated by a disordered phase, as was also pointed out in Sec. XII.4. The only region where a direct transition seem to appear is in the $K_{+}\to 0$ limit, but in this case, the static-order phase is supposed to be unstable due to droplets as discussed in Sec. IX, and therefore this direct transition is only observed at finite systems.
Fig. 42: Additional cuts through the 3D phase diagram in the general asymmetric case. (a) $R$ and (b) $\mathcal{L}$ color maps as a function of $\tilde{K}_{+}$ and $\tilde{K}_{-}$ with $\tilde{J}=1.5$ and $L=80$. (c) $R$ and (d) $\mathcal{L}$ color maps as a function $\tilde{K}_{+}$ and $\tilde{J}$ with $\tilde{K}_{-}=0.3$ and $L=80$. Obtained from MC simulations.
Appendix G Supplementary movies
Movie 1: Evolution in 2D from droplet-induced oscillations to disorder caused by spiral defects. Obtained from MC simulations with parameters: $L=150$, $\tilde{J}=2.8$, and $\tilde{K}=0.3$.
Movie 2: 3D stable swap phase at steady state. Obtained from MC simulations with parameters: $L=80$, $\tilde{J}=1.5$, and $\tilde{K}=0.1$.
Movie 3: 3D system coarsening from a fully disordered state to scroll waves, and finally to global oscillations. Obtained from MC simulations with parameters: $L=80$, $\tilde{J}=2.1$, and $\tilde{K}=0.3$. Line defects of scroll waves are found by coarse-graining the microscopic system as explained in Sec. X with $l=3$ and $s_{\rm max}=0.2$.
Movie 4: Destruction of the static-order phase in 2D by droplet growth, shown as a function of system size. Obtained from MC simulations with parameters: $\tilde{J}=2.8$ and $\tilde{K}=0.3$.
Movie 5: Droplet-capture mechanism in the case of general asymmetric interaction in a 2D system. Obtained from MC simulations with parameters: $L=150$, $\tilde{J}=3.3$, $\tilde{K}_{+}=0.1$, and $\tilde{K}_{+}=0.3$.